### v1

# A Practical Approach for Exploring Granger Connectivity in High-Dimensional Networks of Time Series

Abstract
This paper presents a novel method for discovering effective connectivity between specified pairs of nodes in a high-dimensional network of time series. To accurately perform Granger causality analysis from the first node to the second node, it is essential to eliminate the influence of all other nodes within the network. The approach proposed is to create a low-dimensional representation of all other nodes in the network using frequency-domain-based dynamic principal component analysis (spectral DPCA). The resulting scores are subsequently removed from the first and second nodes of interest, thus eliminating the confounding effect of other nodes within the high-dimensional network. To conduct hypothesis testing on Granger causality, we propose a permutation-based causality test. This test enhances the accuracy of our findings when the error structures are non-Gaussian. The approach has been validated in extensive simulation studies, which demonstrate the efficacy of the methodology as a tool for causality analysis in complex time series networks. The proposed methodology has also been demonstrated to be both expedient and viable on real datasets, with particular success observed on multichannel EEG networks.
1 Introduction
Identifying Granger-causal connections between nodes in high-dimensional networks in neuroscience and network science has significant outcomes. This is particularly relevant when analyzing EEG/fMRI data, where the attention is on a few channels/regions of interest within an extensive network of interacting nodes (Friston, 2011; Friston et al., 2013; Yang et al., 2016; Tu et al., 2019). Conventional Granger causality (GC) is the standard approach to infer causal relationships in high-dimensional EEG networks. However, they are subject to substantial limitations due to the impact of extraneous channels on the Channels of Interest (COI). Accordingly, there is a need for an effectual and parsimonious method to isolate and accurately evaluate the COI’s causal interactions, especially in the presence of multiple channels. The prevailing approaches to causality analysis in high-dimensional networks typically depend on F-tests for Granger causality, which are predicated on the assumption of Gaussianity. However, real-world datasets frequently violate this assumption, leading to biased and unreliable results. Moreover, these methodologies need improvement in effectively removing the influence of non-involved nodes on the COI, which would otherwise lead to misrepresenting causal relationships. On the other hand, conventional dimension reduction techniques in the time domain, such as principal component analysis (PCA), cannot fully capture the data’s frequency-specific characteristics, which are crucial for understanding neural dynamics.
We propose a practical and easily accessible approach based on spectral dynamic principal component analysis (sDPCA) in the frequency domain, combined with permutation testing for Granger causality. The method sDPCA, initially proposed by Brillinger (1969, 1981), extends static PCA by incorporating temporal dependency within the data, thereby reducing the dimensionality of time series in the frequency domain while retaining maximal information. Various methodologies have evolved from sDPCA. For instance, Stoffer (1999) developed a spectral envelope using sDPCA to identify common signals across multiple time series, while Hörmann et al. (2015) proposed a functional variant of sDPCA for analyzing functional time-series data. Ombao and Ho. (2006), Yuxiao Wang et al. (2019), and Ombao and Pinto (2022) indicated that spectral principal component analysis (sPCA) in the frequency domain offers practical benefits and is highly effective in low-dimensional representations of high-dimensional time series. The main objective of this study is to illustrate how the proposed approach enables the practical and sufficient isolation of COI by partialling out the influence of disinterested nodes, which are derived as a dynamic summary in a low-dimensional representation. The frequency-domain dynamic PCA method maintains the spectral characteristics of the data, which are crucial for accurately determining causal relationships in time series networks. By employing a permutation-based GC test in lieu of the conventional F-test, we can relax with the assumption of normality in the error terms, thereby assuring robustness to deviations from normality and rendering the proposed approach more suitable for real-world data.
The proposed approach offers two main advantages. Firstly, DPCA in the frequency domain offers an accurate and practical dimension reduction while preserving the spectral information. This is also essential because EEGs are correlated across time, and hence, the spectrum captures cross-covariances at all lags. This leads to a more precise isolation of COI, reducing the likelihood of spurious causal inferences. Secondly, the permutation-based GC test improves the reliability of causality detection by not relying on the assumption of Gaussianity. This allows the test to accommodate distributions that deviate significantly from normality, such as the fat-tailed distributions observed often in EEG data. The procedure is straightforward and practical to apply. The subsequent sections will elaborate on the proposed approach and its components. Section 2 introduces the Spectral DPCA method and the permutation GC test. Section 3 outlines the scope of the simulation study and presents related results. Section 4 applies our approach to real EEG data and discusses the findings.
2 Methodology
2.1 Inference in high-dimensional networks of time series
Define $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ to be a high dimensional network of time series with possibly $n\geq 256$ nodes. As we consider in this paper, this network may represent the activity of a brain through EEG recordings in which each component (i.e., node or channel) is actually a time series. In this notation of network $N$, $Y^{*}$ and $X^{*}$ denote any two nodes of interest. For problems whose solutions are sought in component-wise, we set our goal as exploring the interaction or efficient connectivity between any pair of nodes in a network rather than the relationality of the whole network. These two nodes are subject to the influence of other nodes due to the network’s relational structure. This means that to dissect the respective efficient connection between any two nodes of interest, they need to be freed from the influences of other nodes. Thereby, the interference-free counterparts of the nodes of interest can be obtained in the following manner:
${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}Y_{t}=Y^{*}_{t}-E\big{(}Y^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}X_{t}=X^{*}_{t}-E\big{(}X^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
where the $f_{z}(.)$ is a function that ensures the dynamic variability of excluded nodes. Practically, the function $f_{z}(.)$ acts as a filter and can take any form as long as it provides the net summary discharge of the nodes excluded. Within the scope of this study, we utilize the spectral domain dynamic PCA implementation made available by Hörmann et al. (2015) via R package called freqdom. This method, originally proposed by Brillinger (1981), enables us to obtain principal component scores that concisely encapsulate variability in multi-dimensional time series, specifically with regard to their serial dependence (see, e.g., Shumway and Stoffer (2017), Ombao and Pinto (2022)). Hence, we propose the filter $f_{z}(.)$ over excluded nodes to provide the following summary indicators,
$f_{z}(Z_{1,t},\ldots,Z_{n-2,t})=\{pc1,pc2,\ldots,pc_{k},\text{interactions}\},$
where $pc1,pc2,\ldots,pc_{k}$ are the scores of the principal components, and interactions represent any interaction terms between these principal components.
The conditional means with respect to these indicators are then given by:
$E\big{(}Y^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=a_{1}\cdot pc1+a_{2}\cdot pc2+\cdots+a_{k}\cdot pc_{k}+\sum a_{ij}\cdot(pc_{i}\times pc_{j}),$
$E\big{(}X^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=b_{1}\cdot pc1+b_{2}\cdot pc2+\cdots+b_{k}\cdot pc_{k}+\sum b_{ij}\cdot(pc_{i}\times pc_{j}),$
where $pc_{i}$ represents the score of the $i$-th dynamic principal component and $a_{i},b_{i},a_{ij},b_{ij}$ are coefficients corresponding to the principal components and their interactions. The number of principal components $k$ can be chosen to be relatively low compared to the number of dimensions, provided that they capture a sufficient proportion of the total variability in the network. For practical implementation, the first few principal components are generally sufficient.
Spectral DPCA
The frequency domain dynamic PCA decomposes a multivariate time series into uncorrelated components in the frequency domain, maximizing the long-run variance explained. Unlike classical PCA, which produces components uncorrelated in space, sDPCA focuses on the spectral density of the data, producing components uncorrelated in time.
Assume a multivariate time series $\mathbf{Z}$ of dimension $T\times n$, where $T$ is the number of time points, and $n$ is the number of variables; this can be considered as a network of time series. The first step in sDPCA involves estimating the spectral density matrix $F(\omega)$ from the time series. The cross-spectral density for $\mathbf{Z}$ can be defined as:
$F_{\mathbf{Z}}(\omega)=\sum_{h\in\mathbb{Z}}\text{Cov}(\mathbf{Z}_{h},\mathbf{Z}_{0})\mathrm{exp}{(-2\pi ih\omega)},$
where $\text{Cov}(\mathbf{Z}_{h},\mathbf{Z}_{0})$ is the auto-covariance matrix at lag $h$. The empirical cross-spectral density can be estimated using a windowed version of the lagged auto-covariance matrices:
$\hat{F}_{\mathbf{Z}}(\omega)=\sum_{|h|\leq q}w\left(\frac{|h|}{q}\right)\hat{C}_{\mathbf{Z}}(h)\mathrm{exp}{(-2\pi ih\omega)},$
where $w(\cdot)$ is a window function (i.e., kernel), $q$ is the window size, and $\hat{C}_{\mathbf{Z}}(h)$ is the empirical lagged auto-covariance given by:
$\hat{C}_{\mathbf{Z}}(h)=\frac{1}{T}\sum_{t=1}^{T-|h|}(\mathbf{Z}_{t+|h|}-\bar{\mathbf{Z}})(\mathbf{Z}_{t}-\bar{\mathbf{Z}})^{\prime},$
for $h\geq 0$, and
$\hat{C}_{\mathbf{Z}}(h)=\frac{1}{T}\sum_{t=|h|+1}^{T}(\mathbf{Z}_{t+|h|}-\bar{\mathbf{Z}})(\mathbf{Z}_{t}-\bar{\mathbf{Z}})^{\prime},$
for $h<0$.
Once the spectral density matrix $F(\omega)$ is obtained, the dynamic principal component filters $\phi_{k}^{(j)}$ are computed, where $k\in[-q,q]\subset\mathbb{Z}$ and the index $j$ is referring to the $j$-th largest dynamic eigenvalue. The filters are derived as the Fourier coefficients of the dynamic eigenvectors $\varphi_{j}(\omega)$ of the spectral density matrix $F(\omega)$:
[…]
The dynamic principal component scores $dpc_{t}^{(j)}$ are then calculated by filtering the time series $\mathbf{Z}_{t}$ with the dynamic principal component filters:
$dpc_{t}^{(j)}=\sum_{k=-q}^{q}\phi_{k}^{(j)}\mathbf{Z}_{t-k}.$
This convolution operation can be represented in matrix notation as:
$\mathbf{dpc}=\mathbf{Z}\ast\mathbf{\Phi},$
where $\ast$ denotes the convolution operation.
The proportion of variance explained by the $j$-th dynamic principal component is quantified as:
[…]
where $\lambda_{j}(\omega)$ is the $j$-th dynamic eigenvalue of the spectral density matrix $F(\omega)$. This ratio provides a measure of how much of the total variance in the time series is captured by each dynamic principal component.
In summary, the sDPCA method involves estimating the spectral density matrix from the time series data using windowed empirical lagged auto-covariances and Fourier transform, computing dynamic principal component filters by performing eigendecomposition on the spectral density matrix and applying the inverse Fourier transform to the dynamic eigenvectors, obtaining dynamic principal component scores by filtering the time series with the computed filters, and quantifying the proportion of variance explained by each dynamic principal component. These steps collectively facilitate the decomposition of multivariate time series into components that are uncorrelated in time and maximize the long-run variance explained.
2.2 Permutation-based Granger Causality test
Granger Causality (GC) is a statistical hypothesis test used to determine whether one time series can predict another. The concept, introduced by Granger (1969), is based on the principle that if the prediction of a time series $Y_{t}$ can be improved by incorporating past values of another time series $X_{t}$, then $X_{t}$ is said to "Granger-cause" $Y_{t}$.
The Granger causality test involves the following steps. First, consider two time series $X_{t}$ and $Y_{t}$. To test if $X_{t}$ Granger-causes $Y_{t}$, we compare the following two models:
- Unrestricted Model:
$Y_{t}=\alpha_{0}+\sum_{i=1}^{p}\alpha_{i}Y_{t-i}+\sum_{j=1}^{q}\beta_{j}X_{t-j}+\epsilon_{t}$
- Restricted Model:
$Y_{t}=\gamma_{0}+\sum_{i=1}^{p}\gamma_{i}Y_{t-i}+\eta_{t}$
Here, $p$ and $q$ are the maximum lags considered, $\alpha_{0}$ and $\gamma_{0}$ are constants, $\alpha_{i}$ and $\gamma_{i}$ are coefficients of the lagged $Y_{t}$, $\beta_{j}$ are coefficients of the lagged $X_{t}$, and $\epsilon_{t}$ and $\eta_{t}$ are error terms.
To determine if $X_{t}$ Granger-causes $Y_{t}$, we test the null hypothesis:
$H_{0}:\beta_{1}=\beta_{2}=\cdots=\beta_{q}=0$
against the alternative hypothesis:
$H_{1}:\text{At least one }\beta_{j}\neq 0\text{ for }j=1,2,\ldots,q$
The test statistic is based on the F-distribution and is computed as:
$F=\frac{(RSS_{R}-RSS_{U})/q}{RSS_{U}/(T-p-q-1)},$
where $RSS_{R}$ and $RSS_{U}$ are the residual sum of squares of the restricted and unrestricted models, respectively, and $T$ is the number of observations. If the computed F-statistic is greater than the critical value from the F-distribution, we reject the null hypothesis, concluding that $X_{t}$ Granger-causes $Y_{t}$.
Granger causality is widely used in fields such as economics, neuroscience, and environmental sciences. In neuroscience, it helps to understand the directional interactions between brain regions by analyzing EEG or fMRI data. However, conventional GC assumes that error terms are normally distributed and homoscedastic, which is often violated in practice, especially in EEG time series that exhibit heteroscedasticity and fat-tailed distributions. These violations can bias results.
Unobserved confounders can also complicate GC analysis. If external variables influencing both time series are omitted, the test might falsely identify causal relationships. This issue is particularly acute in complex systems with many interacting variables.
Permutation-based Granger causality (GC) testing offers a robust alternative to conventional GC tests by relaxing the Gaussianity assumption. The permutation-based GC test evaluates the significance of causality by comparing the observed test statistic against a distribution of test statistics generated through random permutations of the data. This approach is more robust to deviations from normality and other potential violations of the assumptions underlying conventional GC.
To perform the permutation test, we first calculate the test statistic, $F$, for the observed data. We then generate a distribution of the test statistic under the null hypothesis by randomly permuting the time series $X_{t}$ and recomputing the F-statistic for each permutation. Specifically, we shuffle $X_{t}$ to break any potential temporal structure and repeat the GC test for each permuted dataset. This creates a distribution of F-statistics that represent the null distribution.
The permutation p-value is then computed as the proportion of permuted test statistics that are greater than or equal to the observed test statistic:
$p_{\text{perm}}=\frac{\sum_{k=1}^{N}\mathbf{1}(F_{k}\geq F_{\text{obs}})}{N}$
where $N$ is the number of permutations, $F_{k}$ is the test statistic for the $k$-th permutation, $F_{\text{obs}}$ is the observed test statistic, and $\mathbf{1}(\cdot)$ is the indicator function. The indicator function $\mathbf{1}(F_{k}\geq F_{\text{obs}})$ takes the value 1 if $F_{k}$ is greater than or equal to $F_{\text{obs}}$ and 0 otherwise. This function effectively counts the number of times the permuted test statistic exceeds the observed test statistic, providing a measure of how extreme the observed value is within the context of the null distribution.
This permutation approach provides a non-parametric method to assess the significance of the causality, making it more flexible and robust compared to traditional methods. By comparing the observed F-statistic to the distribution generated from permuted data, we account for the actual data distribution and potential dependencies, leading to more reliable inference.
3 Simulation study
The aim of the simulation study is to evaluate the effectiveness of spectral dynamic principal component analysis (sDPCA) combined with Granger causality (GC) testing in identifying true causal relationships within a high-dimensional network of time series data, specifically focusing on EEG data. The study investigates how well the proposed method isolates the channels of interest (COI) from the influences of other channels and accurately assesses their causal interactions.
The study begins by generating a set of 10 channels as mixtures of latent processes proposed by Ombao and Pinto (2022). Each latent process is modeled as an autoregressive process of order 2 (AR(2)). Mathematically, the channels and corresponding indexes are represented as:
$Z_{i}(t)=a_{i}\cdot\delta(t)+b_{i}\cdot\theta(t)+c_{i}\cdot\alpha(t)+d_{i}\cdot\beta(t)+e_{i}\cdot\gamma(t)+\epsilon_{i}(t)$
for $i=1,\ldots,5,11,\ldots,15$, where $\delta,\theta,\alpha,\beta,\gamma$ are latent AR(2) processes, and $\epsilon_{i}(t)$ are error terms.
Next, we generate another set of 10 channels, where each channel is caused by a corresponding channel from the first set with added noise. This can be represented as:
$Z_{i+5}(t)=\phi_{1}Z_{i}(t-1)+\phi_{2}Z_{i}(t-2)+\phi_{3}Z_{i}(t-3)+\phi_{4}Z_{i}(t-4)+\epsilon_{i+5}(t)$
for $i=1,\ldots,5,11,\ldots,15$. The error terms $\epsilon_{i+5}(t)$ follow a specified distribution (e.g., t-distribution with degrees of freedom 15, normal distribution, or non-normal distribution with specified skewness and kurtosis).
The ground truth for Granger causality in this setup is defined by the following relationships:
$Z_{1}\to Z_{6},\quad Z_{2}\to Z_{7},\quad Z_{3}\to Z_{8},\quad Z_{4}\to Z_{9},\quad Z_{5}\to Z_{10},\quad Z_{11}\to Z_{16},\quad Z_{12}\to Z_{17},\quad Z_{13}\to Z_{18},\quad Z_{14}\to Z_{19},\quad Z_{15}\to Z_{20}$
The network is then extended by generating an additional 236 channels ($Z_{21}$ to $Z_{256}$) that share common latent processes with the first 20 channels. This ensures connectivity within the network. Some of these new channels are allowed to have linear influences on the first 20 channels and can be defined as:
$Z_{i}(t)=Z_{i}(t)+\sum_{j\in\text{influencing set}}\text{weight}\cdot Z_{j}(t)$
The number of influencing channels and the weight of their influence are varied in different scenarios to study their effect on the performance of the GC tests.
The study then applies sDPCA to the extended network, focusing on the 236 additional channels to isolate the COI (the first 20 channels). The principal component scores are used to remove the effects of the other channels on the COI. The adjusted time series are obtained by regressing the COI on the principal component scores and retaining the residuals. Bivariate component-wise VAR(1) models are fitted to the residuals to compute the F-statistics for the GC tests. These F-statistics are then compared to critical values from the F-distribution to decide on the presence of GC.
Finally, permutation-based F-statistics are computed by repeatedly permuting the predictor time series and recalculating the F-statistics. The proportion of permuted F-statistics that exceed the observed F-statistic provides the permutation p-value, offering a robust measure of causality that accounts for the actual data distribution.
The results are sensitive to the choice of latent processes and the influence of other channels. Different scenarios are explored by varying the number of influencing channels and the weight of their influence, with the goal of assessing the robustness and reliability of the proposed method in detecting true causal relationships within high-dimensional time series networks. For example, Table 1 presents the accuracy results for one of the simulation studies for different numbers of influencing nodes and varying weights on the COI in a high-dimensional network. The COI consists of the first 20 nodes out of a total of 256 nodes.
Influencing Nodes = 10
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.90 0.85 0.60 0.35
6 1.00 0.90 0.65 0.40
12 0.85 0.85 0.60 0.45
24 0.85 0.75 0.70 0.25
Influencing Nodes = 30
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.90 0.35 0.15 0.30
6 0.85 0.80 0.30 0.25
12 1.00 0.85 0.40 0.40
24 0.95 0.60 0.35 0.20
Influencing Nodes = 60
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.80 0.55 0.25 0.20
6 0.70 0.35 0.20 0.20
12 0.95 0.45 0.20 0.45
24 0.90 0.10 0.30 0.20
Influencing Nodes = 90
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.90 0.30 0.20 0.15
6 0.85 0.15 0.30 0.20
12 0.80 0.25 0.30 0.30
24 0.85 0.40 0.35 0.25
Table 1: Accuracy results for varying numbers of influencing nodes and weights on the COI in a high-dimensional network under a different non-normal error mechanism with different skewness and kurtosis parameters, $\epsilon\sim\mathrm{non-Norm}(\mu=0,\sigma=1,\gamma=0.5,\kappa=0.5).$ The COI consists of the first 20 nodes out of a total of 256 nodes. The numbers 10, 30, 60, 90 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, 0.7 represent the weights of this influence.
This simulation evaluates the performance of the sDPCA method with non-normal error terms, adding complexity due to potential spurious causal relationships. The study employs permutation-based Granger causality testing, detailed in Section 2, to robustly evaluate causal relationships without assuming normally distributed errors.
With 10 influencing nodes, sDPCA shows high accuracy, especially at lower weights (0.1), achieving 0.90 accuracy with 3 Ndpc. However, accuracy declines with higher weights (0.5 and 0.7), dropping to 0.60 and 0.35, respectively, indicating challenges in isolating true causal associations under heavier influences.
As the number of influencing nodes increases to 30, a similar pattern emerges. High accuracy persists at lower weights (0.1), but significantly decreases with higher weights, dropping to 0.30 at a weight of 0.7, underscoring the difficulty in maintaining accuracy with greater external influence.
With 60 and 90 influencing nodes, the challenge is more pronounced. For instance, with 60 influencing nodes and a weight of 0.7, accuracy falls to 0.20, highlighting the significant impact of increased external influences on the COI, complicating the isolation and accurate determination of true causal relationships.
Overall, this table shows that while the sDPCA method can effectively isolate the COI and maintain high accuracy in the presence of external influences, its performance is sensitive to the weight of the influence and the number of influencing nodes. Lower weights and fewer influencing nodes generally result in higher accuracy, while higher weights and more influencing nodes pose greater challenges. These findings are consistent with the theoretical expectations discussed in our simulation framework, where the ability to isolate COI and accurately assess causal interactions is influenced by the degree of external interference. The use of permutation-based GC testing ensures that the assessment of causality is robust and not biased by the assumptions of normality, providing a more reliable analysis of the true causal relationships in the network.
Tables for additional simulations using different error mechanisms to generate the time series are provided in the appendix.
4 Application to EEG data
In this section, we demonstrate the utility of the proposed approach. For practical purposes, we focus on one epoch of the data and analyze four channels: AF3, AF4, C3, and C4. The dataset comes from an experiment involving a series of motor/imagery tasks, providing rich insights into the temporal dynamics of brain activity. The original EEG recordings were collected from 109 volunteers using the BCI2000 system with a 64-channel setup, and the data is publicly available at PhysioNet.org (Schalk et al., 2004). The experiment included 14 runs for each subject: two baseline runs (one with eyes open and one with eyes closed) and twelve task-oriented runs. During the task runs, participants either performed or imagined movements in response to visual cues, such as opening and closing fists or feet. The tasks were divided into categories: physical movement of the left or right fist, imagined movement of the left or right fist, physical movement of both fists or both feet, and imagined movement of both fists or both feet.
Our main objective is to observe the effect of the sDPCA method on dimensionality reduction and its impact on the isolated channels. Specifically, we aim to determine whether this method distorts the temporal information in the time series. Consequently, we focus on only a few channels and share the results as an example, rather than conducting a detailed multichannel analysis. This allows us to provide a demonstration of the approach’s effectiveness in isolating channels of interest and assessing their causal interactions.
In Figure 1, the obtained three principal component (PC) scores are plotted over the time series data from which they were derived. This illustrates that the PC scores can capture the temporal structures of the high-dimensional time series from which they originate. This indicates that the obtained scores preserve the main characteristics of the time series, accurately representing their dynamics. This is a necessary finding that emphasizes the effectiveness and reliability of the procedure.
Figure 1: Isolating Nodes via Spectral Dynamic PCA method
Within the scope of this study, we compared the conditions before and after applying the spectral dynamic principal component analysis (sDPCA) method to the channels of interest. As illustrated in Figures 2 to 3, although the temporal characteristics of the series remain unchanged, there is a noticeable trend toward stationarity in the serşal correlation amplitudes, which initially showed signs of non-stationarity. This adjustment indicates that while the sequential properties of the series are preserved, they become more aligned with the assumptions required by the auxiliary model used for Granger causality (GC) testing.
Figure 2: Sample vs Isolated channels C3 and C4.
Figure 3: Sample vs Isolated channels AF3 and AF4.
Moreover, our test results, summarized in Table 1, reveal significant differences in GC tests conducted before and after the isolation process. In several instances, the GC tests showed contrasting results after the isolation, underlining the presence of spurious GC causality in the original analysis. Specifically, we observed that causal relationships detected in the pre-isolation data were not present in the post-isolation data. This suggests that the influences of other channels, when not accounted for, can lead to misleading conclusions about causality.
Direction of GC GC result Before GC result After
C3 $\rightarrow$ AF3 Cause Do not cause
C3 $\rightarrow$ AF4 Cause Cause
C4 $\rightarrow$ AF3 Do not cause Do not cause
C4 $\rightarrow$ AF4 Cause Cause
AF3 $\rightarrow$ C3 Cause Do not cause
AF3 $\rightarrow$ C4 Cause Cause
AF4 $\rightarrow$ C3 Cause Cause
AF4 $\rightarrow$ C4 Do not cause cause
Table 2: Comparison of Granger Causality Before and After Applying sDPCA
Overall, these findings highlight the effectiveness of the sDPCA method in refining the data by removing extraneous influences, thereby providing a more accurate assessment of the true causal relationships within the EEG channels. This approach enhances the reliability of GC tests, ensuring that the detected causality is genuine and not a result of spurious interactions. The findings accentuate the capability of the sDPCA method to reduce the dimensionality of the data without losing temporal information.
5 Discussion and Conclusion
The study aims at improving the accuracy of Granger causal analysis in high-dimensional time-series networks, specifically focusing on EEG datasets. Through the use of the sDPCA method, we have been able to effectively isolate the COI from the influence of other channels within the network. The results of the simulation studies and the analysis of the real EEG signals show the simplicity, usefulness, and effectiveness of the proposed method.
The simulation study illustrates the challenges of identifying true causal interactions due to potential spurious causality, especially given the presence of non-normality. Integrating sDPCA and permutation-based GC testing significantly improved the accuracy of causal inference. The results show that the method maintained high accuracy in the presence of a moderate number of influencing nodes. However, the method’s accuracy decreases as the number of influencing nodes increases or approaches super/over connectivity (i.e., swarm-like neuronal connectivity). These findings underscore the sensitivity of GC to external influences and stress the necessity of accurately modeling such complexities in real-world data.
The application of the proposed approach to a network of EEG signals provided empirical validation of its potential utility. The principal component scores derived from the spectral domain accurately captured the temporal dynamics of the high-dimensional signals. The results indicate a convergence towards stationarity in the serial correlation amplitudes after applying sDPCA, getting the data closer to the assumptions demanded for reliable GC testing. The comparison of pre- and post-isolation GC results reveals the presence of spurious causality identification in the pre-isolation data, underlining the need to remove extraneous influences for accurate causality assessment.
Despite promising findings, there are several limitations to our study. Although comprehensive, the simulation framework may only capture some nuances of real EEG data. In addition, the test results can be significantly affected by the choice of hyper-parameters, such as the number of principal components, the weights, and the number of influencing nodes. These parameters need to be selected according to the specific characteristics of the datasets to be studied. Furthermore, our method’s reliance on linearity-based PCA in the frequency domain assumes that the underlying temporal dynamics can be adequately captured, which may not always be the case. Future research should consider the integration of sDPCA with the perspective of capturing non-linear relationships. To further improve the robustness and accuracy of causal inference in complex networks, other dimensionality reduction techniques and advanced GC methods should be targeted. In addition, the applicability of the methodology could be significantly enhanced by extending it to include more sophisticated models of external influences and by incorporating real-time analysis capabilities.
In conclusion, our proposed approach provides a practical and reliable solution for isolating and analyzing causal associations in high-dimensional time series networks. The ability of the method to accurately identify true causal connections while mitigating the effects of confounding influences makes it a valuable tool for researchers in neuroscience and other fields dealing with complex time series data. The promising results of the study suggest that this approach can significantly improve the reliability of causality analysis, paving the way for more accurate and insightful investigations of the causal dynamics of high-dimensional networks.
Appendix A Appendix / supplemental material
Influencing Nodes = 10
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.95 0.40 0.50 0.35
6 0.95 0.40 0.35 0.30
12 0.90 0.75 0.50 0.70
24 0.90 0.65 0.25 0.75
Influencing Nodes = 30
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.90 0.40 0.40 0.25
6 0.90 0.40 0.25 0.40
12 0.80 0.50 0.45 0.45
24 0.80 0.35 0.45 0.40
Influencing Nodes = 60
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.35 0.25 0.45 0.15
6 0.85 0.30 0.50 0.20
12 0.65 0.15 0.05 0.35
24 0.40 0.50 0.25 0.15
Influencing Nodes = 90
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.80 0.35 0.30 0.45
6 0.55 0.30 0.25 0.15
12 0.40 0.40 0.20 0.30
24 0.30 0.30 0.20 0.20
Table 3: Accuracy results for varying numbers of influencing nodes and weights on the COI in a high-dimensional network under a non-normal error mechanism with different causal structure, $\epsilon\sim\mathrm{non-Norm}(\mu=0,\sigma=1,\gamma=0.5,\kappa=0.5).$ The COI consists of the first 20 nodes out of a total of 256 nodes. The numbers 10, 30, 60, 90 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, 0.7 represent the weights of this influence.
Influencing Nodes = 10
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.95 0.75 0.75 0.45
6 0.85 0.75 0.45 0.05
12 0.95 0.75 0.75 0.60
24 1.00 0.90 0.60 0.50
Influencing Nodes = 30
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.95 0.65 0.20 0.35
6 0.90 0.35 0.50 0.20
12 0.95 0.50 0.30 0.30
24 1.00 0.60 0.15 0.40
Influencing Nodes = 60
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.90 0.20 0.25 0.20
6 0.85 0.50 0.20 0.40
12 0.85 0.40 0.25 0.30
24 0.90 0.15 0.15 0.35
Influencing Nodes = 90
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.75 0.35 0.40 0.30
6 0.90 0.25 0.30 0.15
12 0.95 0.60 0.40 0.20
24 0.90 0.25 0.35 0.10
Table 4: Accuracy results for varying numbers of influencing nodes and weights on the COI in a high-dimensional network under non-normal error-generating mechanisms with different skewness and kurtosis parameters, $\epsilon\sim\mathrm{non-Norm}(\mu=-1,\sigma=1,\gamma=1.5,\kappa=3.5).$. The COI consists of the first 20 nodes out of a total of 256 nodes. The numbers 10, 30, 60, 90 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, 0.7 represent the weights of this influence.
Influencing Nodes = 10
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.95 0.90 0.60 0.45
6 0.90 0.90 0.50 0.45
12 0.90 0.90 0.50 0.20
24 1.00 0.95 0.70 0.55
Influencing Nodes = 30
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.95 0.50 0.30 0.30
6 0.85 0.55 0.30 0.20
12 0.95 0.50 0.20 0.10
24 0.85 0.65 0.45 0.25
Influencing Nodes = 60
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.85 0.35 0.05 0.20
6 0.75 0.30 0.25 0.35
12 0.90 0.35 0.15 0.30
24 0.80 0.55 0.45 0.25
Influencing Nodes = 90
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.90 0.25 0.15 0.35
6 0.85 0.40 0.20 0.25
12 0.90 0.25 0.05 0.05
24 1.00 0.20 0.20 0.25
Table 5: Accuracy results for varying numbers of influencing nodes and weights on the COI in a high-dimensional network with Gaussian error-generating mechanisms. The COI consists of the first 20 nodes out of a total of 256 nodes. The numbers 10, 30, 60, 90 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, 0.7 represent the weights of this influence.
Influencing Nodes = 10
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.95 0.65 0.25 0.35
6 0.90 0.65 0.50 0.45
12 0.95 0.55 0.55 0.25
24 0.90 0.65 0.25 0.30
Influencing Nodes = 30
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.90 0.55 0.20 0.10
6 0.70 0.65 0.30 0.30
12 0.75 0.55 0.50 0.35
24 0.75 0.65 0.40 0.30
Influencing Nodes = 60
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.80 0.50 0.35 0.15
6 0.70 0.40 0.20 0.15
12 0.90 0.45 0.20 0.30
24 0.75 0.55 0.20 0.25
Influencing Nodes = 90
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.55 0.25 0.35 0.30
6 0.55 0.60 0.15 0.25
12 0.65 0.20 0.30 0.25
24 0.45 0.30 0.15 0.20
Table 6: Accuracy results for varying numbers of influencing nodes and weights on the COI in a high-dimensional network with another Gaussian error-generating mechanism. The COI consists of the first 20 nodes out of a total of 256 nodes. The numbers 10, 30, 60, 90 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, 0.7 represent the weights of this influence.
Influencing Nodes = 10
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 1.00 0.60 0.70 0.30
6 0.90 0.40 0.40 0.35
12 0.95 0.15 0.25 0.30
24 1.00 0.40 0.45 0.70
Influencing Nodes = 30
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.85 0.35 0.30 0.30
6 0.70 0.40 0.20 0.40
12 0.75 0.35 0.40 0.25
24 0.80 0.50 0.50 0.40
Influencing Nodes = 60
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.40 0.50 0.35 0.15
6 0.50 0.70 0.30 0.25
12 0.50 0.40 0.25 0.25
24 0.40 0.60 0.20 0.40
Influencing Nodes = 90
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.45 0.25 0.20 0.25
6 0.45 0.25 0.15 0.40
12 0.35 0.55 0.25 0.30
24 0.60 0.30 0.40 0.55
Table 7: Accuracy results for varying numbers of influencing nodes and weights on the COI in a high-dimensional network with another Gaussian error-generating mechanism. The COI consists of the first 20 nodes out of a total of 256 nodes. The numbers 10, 30, 60, 90 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, 0.7 represent the weights of this influence.

### v2

@@
-A Practical Approach for Exploring Granger Connectivity in High-Dimensional Networks of Time Series
+Granger Causality in High-Dimensional Networks of Time Series
 Abstract
-This paper presents a novel method for discovering effective connectivity between specified pairs of nodes in a high-dimensional network of time series.
-To accurately perform Granger causality analysis from the first node to the second node, it is essential to eliminate the influence of all other nodes within the network.
-The approach proposed is to create a low-dimensional representation of all other nodes in the network using frequency-domain-based dynamic principal component analysis (spectral DPCA).
-The resulting scores are subsequently removed from the first and second nodes of interest, thus eliminating the confounding effect of other nodes within the high-dimensional network.
-To conduct hypothesis testing on Granger causality, we propose a permutation-based causality test.
-This test enhances the accuracy of our findings when the error structures are non-Gaussian.
-The approach has been validated in extensive simulation studies, which demonstrate the efficacy of the methodology as a tool for causality analysis in complex time series networks.
-The proposed methodology has also been demonstrated to be both expedient and viable on real datasets, with particular success observed on multichannel EEG networks.
-1 Introduction
-Identifying Granger-causal connections between nodes in high-dimensional networks in neuroscience and network science has significant outcomes.
-This is particularly relevant when analyzing EEG/fMRI data, where the attention is on a few channels/regions of interest within an extensive network of interacting nodes (Friston, 2011; Friston et al., 2013; Yang et al., 2016; Tu et al., 2019).
-Conventional Granger causality (GC) is the standard approach to infer causal relationships in high-dimensional EEG networks.
-However, they are subject to substantial limitations due to the impact of extraneous channels on the Channels of Interest (COI).
-Accordingly, there is a need for an effectual and parsimonious method to isolate and accurately evaluate the COI’s causal interactions, especially in the presence of multiple channels.
-The prevailing approaches to causality analysis in high-dimensional networks typically depend on F-tests for Granger causality, which are predicated on the assumption of Gaussianity.
-However, real-world datasets frequently violate this assumption, leading to biased and unreliable results.
-Moreover, these methodologies need improvement in effectively removing the influence of non-involved nodes on the COI, which would otherwise lead to misrepresenting causal relationships.
-On the other hand, conventional dimension reduction techniques in the time domain, such as principal component analysis (PCA), cannot fully capture the data’s frequency-specific characteristics, which are crucial for understanding neural dynamics.
-We propose a practical and easily accessible approach based on spectral dynamic principal component analysis (sDPCA) in the frequency domain, combined with permutation testing for Granger causality.
-The method sDPCA, initially proposed by Brillinger (1969, 1981), extends static PCA by incorporating temporal dependency within the data, thereby reducing the dimensionality of time series in the frequency domain while retaining maximal information.
+A novel approach is developed for discovering directed connectivity between specified pairs of nodes in a high-dimensional network (HDN) of brain signals.
+To accurately identify causal connectivity for such specified objectives, it is necessary to properly address the influence of all other nodes within the network.
+The proposed procedure herein starts with the estimation of a low-dimensional representation of the other nodes in the network utilizing (frequency-domain-based) spectral dynamic principal component analysis (sDPCA).
+The resulting scores can then be removed from the nodes of interest, thus eliminating the confounding effect of other nodes within the HDN.
+Accordingly, causal interactions can be dissected between nodes that are isolated from the effects of the network.
+Extensive simulations have demonstrated the effectiveness of this approach as a tool for causality analysis in complex time series networks.
+The proposed methodology has also been shown to be applicable to multichannel EEG networks.
+Index Terms:
+High-Dimensional Networks of Signals, Connectivity, Causality, Spectral Dynamic PCA, EEG signals
+I Introduction
+Identifying effective connectivity in high-dimensional networks (HDN) is a challenging problem that remains an active area of research and has been previously addressed by several studies [see, e.g., 1, 2, 3, 4, 5, 6, 7, 8].
+Moreover, it is difficult to infer or interpret interactions between a pair of nodes while taking into account potential interactions with all other nodes in the background [9, 10].
+This is a particularly relevant issue (e.g., analysis of EEG-like signals) when the focus might stand on a few channels/nodes of interest (COI/NOI) within an extensive network of interacting nodes.
+In fact, utilizing conventional vector autoregressive models (VAR) can be a practical way to identify pairwise causal interactions in high-dimensional networks as long as the potential confounding effect due to the high-dimensional background nodes is addressed.
+In this paper, we propose a framework that provides a low-dimensional summary of the background activity in the high-dimensional network (HDN).
+To discern effective connectivity, particularly between pairs of nodes in an HDN of brain signals, we propose a practical and easily accessible approach based on spectral dynamic principal component analysis (PCA) ([11, 12]) combined with testing for GC in lieu of the conventional F-test.
+The sDPCA method enhances the classical PCA by incorporating temporal dependency within the data, thereby reducing the dimensionality of time series in the frequency domain while retaining maximal temporal information.
+In other words, while the classical PCA is based on the eigendecomposition of the zero-lag (or contemporaneous) covariance matrix, sDPCA is based on the eigendecomposition of the spectral matrix at all frequencies.
+Since there is a 1-1 relationship between the spectral matrix (at all frequencies) and the covariance matrix (at all lags) then this approach captures lead-lag relationships amongst signals.
 Various methodologies have evolved from sDPCA.
-For instance, Stoffer (1999) developed a spectral envelope using sDPCA to identify common signals across multiple time series, while Hörmann et al.
-(2015) proposed a functional variant of sDPCA for analyzing functional time-series data.
-Ombao and Ho.
-(2006), Yuxiao Wang et al.
-(2019), and Ombao and Pinto (2022) indicated that spectral principal component analysis (sPCA) in the frequency domain offers practical benefits and is highly effective in low-dimensional representations of high-dimensional time series.
+For instance, [13] developed a spectral envelope using sDPCA to identify common signals across multiple time series, while [14] proposed a functional variant of sDPCA for analyzing functional time-series data.
+As indicated by [15, 5, 16], spectral principal component analysis (sPCA) in the frequency domain offers practical benefits and is highly effective in low-dimensional representations of high-dimensional time series.
 The main objective of this study is to illustrate how the proposed approach enables the practical and sufficient isolation of COI by partialling out the influence of disinterested nodes, which are derived as a dynamic summary in a low-dimensional representation.
 The frequency-domain dynamic PCA method maintains the spectral characteristics of the data, which are crucial for accurately determining causal relationships in time series networks.
-By employing a permutation-based GC test in lieu of the conventional F-test, we can relax with the assumption of normality in the error terms, thereby assuring robustness to deviations from normality and rendering the proposed approach more suitable for real-world data.
 The proposed approach offers two main advantages.
-Firstly, DPCA in the frequency domain offers an accurate and practical dimension reduction while preserving the spectral information.
-This is also essential because EEGs are correlated across time, and hence, the spectrum captures cross-covariances at all lags.
+Firstly, sDPCA in the frequency domain offers an accurate and practical dimension reduction while preserving the spectral information.
+This is essential because the spectrum of a signal captures cross-covariances at all lags.
 This leads to a more precise isolation of COI, reducing the likelihood of spurious causal inferences.
-Secondly, the permutation-based GC test improves the reliability of causality detection by not relying on the assumption of Gaussianity.
-This allows the test to accommodate distributions that deviate significantly from normality, such as the fat-tailed distributions observed often in EEG data.
-The procedure is straightforward and practical to apply.
+Secondly, the proposed approach enables the formation of a comprehensive yet interpretable representation of the interactions between specified nodes at a defined level of resolution rather than being unable to infer by the expansive number of possible interactions between nodes.
+This facilitates the study of considerable numbers of interactions, ranging from pairs of nodes to interactions between a tractable number of clusters of nodes.
 The subsequent sections will elaborate on the proposed approach and its components.
-Section 2 introduces the Spectral DPCA method and the permutation GC test.
+Section 2 presents the proposed approach.
 Section 3 outlines the scope of the simulation study and presents related results.
-Section 4 applies our approach to real EEG data and discusses the findings.
-2 Methodology
-2.1 Inference in high-dimensional networks of time series
-Define $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ to be a high dimensional network of time series with possibly $n\geq 256$ nodes.
+Section 4 applies the proposed approach to EEG data and discusses the findings.
+II Methodology
+II-A Inference in high-dimensional networks of time series
+Define $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ to be a high dimensional network of signals with possibly $n\geq 256$ nodes.
 As we consider in this paper, this network may represent the activity of a brain through EEG recordings in which each component (i.e., node or channel) is actually a time series.
 In this notation of network $N$, $Y^{*}$ and $X^{*}$ denote any two nodes of interest.
-For problems whose solutions are sought in component-wise, we set our goal as exploring the interaction or efficient connectivity between any pair of nodes in a network rather than the relationality of the whole network.
+For problems whose solutions are sought component-wise, we set our goal as exploring the interaction or efficient connectivity between any pair of nodes in a network rather than the relationality of the whole network.
 These two nodes are subject to the influence of other nodes due to the network’s relational structure.
-This means that to dissect the respective efficient connection between any two nodes of interest, they need to be freed from the influences of other nodes.
+This means that to dissect the respective efficient connection between any two nodes of interest; they need to be freed from the influences of other nodes.
 Thereby, the interference-free counterparts of the nodes of interest can be obtained in the following manner:
-${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}Y_{t}=Y^{*}_{t}-E\big{(}Y^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
-${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}X_{t}=X^{*}_{t}-E\big{(}X^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
-where the $f_{z}(.)$ is a function that ensures the dynamic variability of excluded nodes.
+$Y_{t}=Y^{*}_{t}-E\big{(}Y^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
+$X_{t}=X^{*}_{t}-E\big{(}X^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
+where the $f_{z}(.)$ is a function that encompasses the dynamic variability of excluded nodes.
 Practically, the function $f_{z}(.)$ acts as a filter and can take any form as long as it provides the net summary discharge of the nodes excluded.
@@
 (2015) via R package called freqdom.
-This method, originally proposed by Brillinger (1981), enables us to obtain principal component scores that concisely encapsulate variability in multi-dimensional time series, specifically with regard to their serial dependence (see, e.g., Shumway and Stoffer (2017), Ombao and Pinto (2022)).
+This method, originally proposed by Brillinger (1981), enables us to obtain principal component scores that concisely encapsulate variability in multi-dimensional time series, specifically with regard to their serial dependence [17, 16, see, e.g., ].
 Hence, we propose the filter $f_{z}(.)$ over excluded nodes to provide the following summary indicators,
 $f_{z}(Z_{1,t},\ldots,Z_{n-2,t})=\{pc1,pc2,\ldots,pc_{k},\text{interactions}\},$
-where $pc1,pc2,\ldots,pc_{k}$ are the scores of the principal components, and interactions represent any interaction terms between these principal components.
+where $pc1,pc2,\ldots,pc_{k}$ are the scores of the principal components obtained through sDPCA, and interactions represent any interaction terms between these principal components.
 The conditional means with respect to these indicators are then given by:
-$E\big{(}Y^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=a_{1}\cdot pc1+a_{2}\cdot pc2+\cdots+a_{k}\cdot pc_{k}+\sum a_{ij}\cdot(pc_{i}\times pc_{j}),$
-$E\big{(}X^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=b_{1}\cdot pc1+b_{2}\cdot pc2+\cdots+b_{k}\cdot pc_{k}+\sum b_{ij}\cdot(pc_{i}\times pc_{j}),$
+$E\big{(}Y^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=a_{1}\cdot pc1+a_{2}\cdot pc+$
+$\cdots+a_{k}\cdot pc_{k}+\sum a_{ij}\cdot(pc_{i}\times pc_{j}),$
+$E\big{(}X^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=b_{1}\cdot pc1+b_{2}\cdot pc2+$
+$\cdots+b_{k}\cdot pc_{k}+\sum b_{ij}\cdot(pc_{i}\times pc_{j}),$
 where $pc_{i}$ represents the score of the $i$-th dynamic principal component and $a_{i},b_{i},a_{ij},b_{ij}$ are coefficients corresponding to the principal components and their interactions.
@@
 for $h\geq 0$, and
-$\hat{C}_{\mathbf{Z}}(h)=\frac{1}{T}\sum_{t=|h|+1}^{T}(\mathbf{Z}_{t+|h|}-\bar{\mathbf{Z}})(\mathbf{Z}_{t}-\bar{\mathbf{Z}})^{\prime},$
+$\hat{C}_{\mathbf{Z}}(h)=\hat{C}_{\mathbf{Z}}(-h)^{\prime}=\frac{1}{T}\sum_{t=1}^{T+h}(\mathbf{Z}_{t}-\bar{\mathbf{Z}})(\mathbf{Z}_{t-h}-\bar{\mathbf{Z}})^{\prime}.$
 for $h<0$.
@@
 These steps collectively facilitate the decomposition of multivariate time series into components that are uncorrelated in time and maximize the long-run variance explained.
-2.2 Permutation-based Granger Causality test
+Terms and Notation
+  • $\mathbf{Z}$: Original multivariate time series matrix of dimensions $T\times n$, where $T$ is the number of time points and $n$ is the number of variables.
+  • $F_{\mathbf{Z}}(\omega)$: Cross-spectral density matrix at frequency $\omega$.
+  • $\text{Cov}(\mathbf{Z}_{h},\mathbf{Z}_{0})$: Covariance matrix at lag $h$.
+  • $\hat{C}_{\mathbf{Z}}(h)$: Empirical lagged auto-covariance matrix at lag $h$.
+  • $w(\cdot)$: Window function used to smooth the estimates.
+  • $q$: Window size for the kernel estimator.
+  • $\phi_{k}^{(j)}$: Dynamic principal component filters, obtained as the Fourier coefficients of the dynamic eigenvectors.
+  • $\varphi_{j}(\omega)$: Dynamic eigenvectors of the spectral density matrix $F(\omega)$.
+  • $dpc_{t}^{(j)}$: Dynamic principal component scores.
+  • $\mathbf{dpc}$: Matrix of dynamic principal component scores.
+  • $\mathbf{\Phi}$: Matrix of dynamic principal component filters.
+  • $\lambda_{j}(\omega)$: $j$-th dynamic eigenvalue of the spectral density matrix $F(\omega)$.
+II-B Granger Causality test
 Granger Causality (GC) is a statistical hypothesis test used to determine whether one time series can predict another.
-The concept, introduced by Granger (1969), is based on the principle that if the prediction of a time series $Y_{t}$ can be improved by incorporating past values of another time series $X_{t}$, then $X_{t}$ is said to "Granger-cause" $Y_{t}$.
+The concept, introduced by Granger (1969), is based on the principle that if the prediction of a time series $Y_{t}$ can be improved by incorporating past values of another time series $X_{t}$, then $X_{t}$ is said to ”Granger-cause” $Y_{t}$.
 The Granger causality test involves the following steps. First, consider two time series $X_{t}$ and $Y_{t}$. To test if $X_{t}$ Granger-causes $Y_{t}$, we compare the following two models:
@@
 This issue is particularly acute in complex systems with many interacting variables.
-Permutation-based Granger causality (GC) testing offers a robust alternative to conventional GC tests by relaxing the Gaussianity assumption.
-The permutation-based GC test evaluates the significance of causality by comparing the observed test statistic against a distribution of test statistics generated through random permutations of the data.
-This approach is more robust to deviations from normality and other potential violations of the assumptions underlying conventional GC.
-To perform the permutation test, we first calculate the test statistic, $F$, for the observed data.
-We then generate a distribution of the test statistic under the null hypothesis by randomly permuting the time series $X_{t}$ and recomputing the F-statistic for each permutation.
-Specifically, we shuffle $X_{t}$ to break any potential temporal structure and repeat the GC test for each permuted dataset.
-This creates a distribution of F-statistics that represent the null distribution.
-The permutation p-value is then computed as the proportion of permuted test statistics that are greater than or equal to the observed test statistic:
-$p_{\text{perm}}=\frac{\sum_{k=1}^{N}\mathbf{1}(F_{k}\geq F_{\text{obs}})}{N}$
-where $N$ is the number of permutations, $F_{k}$ is the test statistic for the $k$-th permutation, $F_{\text{obs}}$ is the observed test statistic, and $\mathbf{1}(\cdot)$ is the indicator function.
-The indicator function $\mathbf{1}(F_{k}\geq F_{\text{obs}})$ takes the value 1 if $F_{k}$ is greater than or equal to $F_{\text{obs}}$ and 0 otherwise.
-This function effectively counts the number of times the permuted test statistic exceeds the observed test statistic, providing a measure of how extreme the observed value is within the context of the null distribution.
-This permutation approach provides a non-parametric method to assess the significance of the causality, making it more flexible and robust compared to traditional methods.
-By comparing the observed F-statistic to the distribution generated from permuted data, we account for the actual data distribution and potential dependencies, leading to more reliable inference.
-3 Simulation study
-The aim of the simulation study is to evaluate the effectiveness of spectral dynamic principal component analysis (sDPCA) combined with Granger causality (GC) testing in identifying true causal relationships within a high-dimensional network of time series data, specifically focusing on EEG data.
+II-C Proposed Approach
+In high-dimensional networks of time series, analyzing causal connectivity between specific pairs of signals can be both challenging and computationally intensive due to the complexity of modeling interactions across all nodes.
+The approach to be followed would be to model the whole material if there were no high dimensionality problems.
+However, modeling under high dimensionality requires fitting highly parameterized models to the entire network, leading to difficult-to-interpret estimates and an increased risk of overfitting.
+To overcome these challenges, we propose a practical and easily implementable method that focuses on isolating the nodes/channels of interest by effectively reducing the dimensionality of the network.
+Our approach utilizes spectral dynamic principal component analysis to capture the essential variability of the background network in a low-dimensional representation.
+By partialling out the influence of other nodes through sDPCA, we can isolate the COI and accurately assess the causal connectivity between them using conventional Granger causality tests.
+This method simplifies the analysis by avoiding the need to model the entire high-dimensional network and enhances interpretability when the primary interest is in the causal relationship between specific signals.
+The proposed approach is given in the Procedure outline below.
+Input: High-dimensional Networks of Signals $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$
+Output: Determination of causal relationship between $Y^{*}_{t}$ and $X^{*}_{t}$
+1. Identify Nodes of Interest: NOI:$\{Y^{*}_{t},X^{*}_{t}\}$.
+In practice, each node can also be a cluster of nodes that can be represented by sDPCA scores.
+2. Apply sDPCA:
+  1. i
+    Estimate spectral density matrix $F_{\mathbf{Z}}(\omega)$ for $\mathbf{Z}_{t}=\{Z_{1,t},\ldots,Z_{n-2,t}\}$.
+  2. ii
+    Compute dynamic eigenvectors $\varphi_{j}(\omega)$ and
+eigenvalues $\lambda_{j}(\omega)$.
+  • iii
+    Derive dynamic principal component filters:
+[…]
+  • iv
+    Calculate dynamic principal component scores:
+    $\text{dpc}_{t}^{(j)}=\sum_{k=-q}^{q}\phi_{k}^{(j)}\mathbf{Z}_{t-k}$
+  • 3. Partial Out Influence of Other Nodes:
+    $Y_{t}=Y^{*}_{t}-\hat{E}\big{(}Y^{*}_{t}\mid\text{dpc}_{t}^{(1)},\ldots,\text{dpc}_{t}^{(k)}\big{)}$
+    $X_{t}=X^{*}_{t}-\hat{E}\big{(}X^{*}_{t}\mid\text{dpc}_{t}^{(1)},\ldots,\text{dpc}_{t}^{(k)}\big{)}$
+    4. Fit Restricted/Unrestricted Models:
+    Unrestricted Model:
+    $Y_{t}=\alpha_{0}+\sum_{i=1}^{p}\alpha_{i}Y_{t-i}+\sum_{j=1}^{q}\beta_{j}X_{t-j}+\epsilon_{t}$
+    Restricted Model:
+    $Y_{t}=\gamma_{0}+\sum_{i=1}^{p}\gamma_{i}Y_{t-i}+\eta_{t}$
+    5. Perform Granger Causality Tests:
+    1. i
+      Test null hypothesis $H_{0}:\beta_{j}=0$
+    for all $j$ in the equation for $Y_{t}$.
+  • ii
+    Compute F-statistic:
+    $F=\frac{(RSS_{R}-RSS_{U})/q}{RSS_{U}/(T-K)}$
+    where:
+    • $RSS_{R}$: Residual sum of squares of
+  • restricted model (without $X_{t-j}$ terms).
+  • $RSS_{U}$: Residual sum of squares of
+  • unrestricted model.
+  • $q$: Number of restrictions (parameters set to zero). $K$: Total number of parameters in the unrestricted model. iii Compare F-statistic with critical value
+  • from $F$-distribution $F_{q,T-K}$.
+  • iv
+    Reject $H_{0}$ if $F$ exceeds critical value; conclude $X_{t}$ Granger-causes $Y_{t}$. v Repeat steps for testing causality from $Y_{t}$ to $X_{t}$.
+  • Procedure 1 Identifying Causal Connectivity in HDN
+    III Simulation study
+    The aim of this simulation study is to evaluate the effectiveness of spectral dynamic principal component analysis (sDPCA) combined with Granger causality (GC) testing in identifying true causal relationships within a high-dimensional network of time series data, specifically focusing on EEG-like data.
 The study investigates how well the proposed method isolates the channels of interest (COI) from the influences of other channels and accurately assesses their causal interactions.
-The study begins by generating a set of 10 channels as mixtures of latent processes proposed by Ombao and Pinto (2022).
-Each latent process is modeled as an autoregressive process of order 2 (AR(2)).
-Mathematically, the channels and corresponding indexes are represented as:
-$Z_{i}(t)=a_{i}\cdot\delta(t)+b_{i}\cdot\theta(t)+c_{i}\cdot\alpha(t)+d_{i}\cdot\beta(t)+e_{i}\cdot\gamma(t)+\epsilon_{i}(t)$
-for $i=1,\ldots,5,11,\ldots,15$, where $\delta,\theta,\alpha,\beta,\gamma$ are latent AR(2) processes, and $\epsilon_{i}(t)$ are error terms.
-Next, we generate another set of 10 channels, where each channel is caused by a corresponding channel from the first set with added noise. This can be represented as:
-$Z_{i+5}(t)=\phi_{1}Z_{i}(t-1)+\phi_{2}Z_{i}(t-2)+\phi_{3}Z_{i}(t-3)+\phi_{4}Z_{i}(t-4)+\epsilon_{i+5}(t)$
-for $i=1,\ldots,5,11,\ldots,15$.
-The error terms $\epsilon_{i+5}(t)$ follow a specified distribution (e.g., t-distribution with degrees of freedom 15, normal distribution, or non-normal distribution with specified skewness and kurtosis).
-The ground truth for Granger causality in this setup is defined by the following relationships:
-$Z_{1}\to Z_{6},\quad Z_{2}\to Z_{7},\quad Z_{3}\to Z_{8},\quad Z_{4}\to Z_{9},\quad Z_{5}\to Z_{10},\quad Z_{11}\to Z_{16},\quad Z_{12}\to Z_{17},\quad Z_{13}\to Z_{18},\quad Z_{14}\to Z_{19},\quad Z_{15}\to Z_{20}$
-The network is then extended by generating an additional 236 channels ($Z_{21}$ to $Z_{256}$) that share common latent processes with the first 20 channels.
-This ensures connectivity within the network.
-Some of these new channels are allowed to have linear influences on the first 20 channels and can be defined as:
-$Z_{i}(t)=Z_{i}(t)+\sum_{j\in\text{influencing set}}\text{weight}\cdot Z_{j}(t)$
-The number of influencing channels and the weight of their influence are varied in different scenarios to study their effect on the performance of the GC tests.
-The study then applies sDPCA to the extended network, focusing on the 236 additional channels to isolate the COI (the first 20 channels).
-The principal component scores are used to remove the effects of the other channels on the COI.
-The adjusted time series are obtained by regressing the COI on the principal component scores and retaining the residuals.
-Bivariate component-wise VAR(1) models are fitted to the residuals to compute the F-statistics for the GC tests.
-These F-statistics are then compared to critical values from the F-distribution to decide on the presence of GC.
-Finally, permutation-based F-statistics are computed by repeatedly permuting the predictor time series and recalculating the F-statistics.
-The proportion of permuted F-statistics that exceed the observed F-statistic provides the permutation p-value, offering a robust measure of causality that accounts for the actual data distribution.
+We conduct two sets of simulations: one with linear influences and another with nonlinear influences.
+    III-A Linear Scheme
+    In the linear scheme of the simulation study, we generate a network of time series data with predefined causal relationships and linear influences from additional channels.
+    III-A1 Generating Initial Channels
+    We begin by generating $N=40$ initial time series channels, denoted by $\{Z_{i}(t)\}_{i=1}^{N}$, over time points $t=1,\ldots,T$.
+The channels are divided into $N/2=20$ causal pairs, where each pair consists of a ”cause” and an ”effect” channel.
+    For each cause channel $Z_{i}(t)$, where $i=1,\ldots,N/2$, we model it as a weighted sum of latent processes representing different EEG frequency bands plus noise:
+    $Z_{i}(t)=a_{i}\,\delta(t)+b_{i}\,\theta(t)+c_{i}\,\alpha(t)+d_{i}\,\beta(t)+e_{i}\,\gamma(t)+\varepsilon_{i}(t),$
+    where:
+    • $\delta(t)$, $\theta(t)$, $\alpha(t)$, $\beta(t)$, and $\gamma(t)$ are latent processes corresponding to the EEG frequency bands:
+      • Delta (0.50.50.50.54444 Hz): $\delta(t)$,
+      • Theta (44448888 Hz): $\theta(t)$,
+      • Alpha (888812121212 Hz): $\alpha(t)$,
+      • Beta (1212121230303030 Hz): $\beta(t)$,
+      • Gamma ($>$ $30$ Hz): $\gamma(t)$.
+      Each latent process is modeled as an autoregressive process of order 2 (AR(2)):
+      $\nu(t)=\phi_{1\nu}\nu(t-1)+\phi_{2\nu}\nu(t-2)+\eta_{\nu}(t),\quad\nu\in\{\delta,\theta,\alpha,\beta,\gamma\},$
+      where $\eta_{\nu}(t)$ is white noise. The AR coefficients can expressed in terms of $r_{\nu}$ and $\varrho_{\nu}$:
+      $\phi_{1\nu}=2r_{\nu}\cos(\varrho_{\nu}),$
+      $\phi_{2\nu}=-r_{\nu}^{2}.$
+      where $r_{\nu}$ is the damping factor, chosen close to $1$ (e.g., $10.8\leq r_{\nu}<$ 1) to ensure stationarity and control the bandwidth of the spectral peak.
+And, $\varrho_{\nu}$ is the angular frequency (in radians per sample), related to the desired frequency $f_{\nu}$ (in Hz) by:
+      $\varrho_{\nu}=2\pi\frac{f_{\nu}}{f_{s}},$
+      with $f_{s}$ being the sampling frequency. By modeling the latent processes to correspond to specific EEG frequency bands, we aim to simulate EEG-like signals more realistically.
+    • $a_{i}$, $b_{i}$, $c_{i}$, $d_{i}$, and $e_{i}$ are random weights drawn from a uniform distribution to ensure variability among channels.
+    • $\varepsilon_{i}(t)$ is white noise added to each channel.
+    For each effect channel $Z_{j}(t)$, where $j=i+N/2$ and $i=1,\ldots,N/2$, we generate it as an autoregressive process influenced by its corresponding cause channel:
+    $Z_{j}(t)=\sum_{k=1}^{p}\phi_{k}Z_{i}(t-k)+\varepsilon_{j}(t),$
+    where:
+    • $p=2$ is the order of the autoregressive process.
+    • $\phi_{k}$ are AR coefficients ensuring the stability of the process.
+    • $\varepsilon_{j}(t)$ is white noise added to the effect channel.
+    III-A2 Extending the Network
+    We extend the network by generating additional $N_{\text{ext}}=472$ channels, resulting in a total of $N_{\text{expanded}}=512$ channels.
+These additional channels $\{Z_{i}(t)\}_{i=N+1}^{N_{\text{expanded}}}$ are generated similarly to the cause channels but without being part of the predefined causal pairs:
+    $Z_{i}(t)=a_{i}\,\delta(t)+b_{i}\,\theta(t)+c_{i}\,\alpha(t)+d_{i}\,\beta(t)+e_{i}\,\gamma(t)+\varepsilon_{i}(t).$
+    III-A3 Introducing Linear Influences
+    To simulate the effect of a high-dimensional network on the initial channels, we introduce linear influences from a subset of the additional channels.
+We select a set of connected nodes $\mathcal{C}\subset\{N+1,\ldots,N_{\text{expanded}}\}$ of size nofcn (number of connected nodes).
+    For each initial channel $Z_{i}(t)$, $i=1,\ldots,N$, we randomly select $k$ influencing nodes from $\mathcal{C}$ and assign influence weights $W_{ij}$ drawn from a uniform distribution.
+    The linear influence on channel $i$ at time $t$ is defined as:
+    $L_{i}(t)=\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t),$
+    where $\mathcal{I}_{i}$ is the set of influencing nodes for channel $i$.
+For each initial channel $Z_{i}(t)$ and each influencing node $Z_{j}(t)\in\mathcal{I}_{i}$, we assign an influence weight $W_{ij}$ drawn from a uniform distribution centered around a specified influence weight $\omega$:
+    $W_{ij}\sim\text{Uniform}\left(\omega-\frac{\omega}{16},\,\omega+\frac{\omega}{16}\right),\quad\text{for }j\in\mathcal{I}_{i}.$
+    We then update the initial channels by adding the linear influence:
+    $\tilde{Z}_{i}(t)=Z_{i}(t)+L_{i}(t).$
+    The updated initial channels $\tilde{Z}_{i}(t)$ now contain both their original dynamics and the linear influence from the connected nodes.
+    III-B Nonlinear Scheme
+    In the nonlinear scheme, we introduce moderate nonlinearity into the influence of the connected nodes on the initial channels to study the robustness of the sDPCA and GC testing framework under nonlinear conditions.
+    III-B1 Introducing Nonlinear Influences
+    Similar to the linear simulation, we select a set of connected nodes $\mathcal{C}$ and assign influence weights $W_{ij}$ for each initial channel $Z_{i}(t)$.
+    The linear influence is computed as before:
+    $L_{i}(t)=\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t).$
+    To introduce nonlinearity, we compute interaction terms between the influencing nodes for each initial channel:
+    $Q_{i}(t)=\sum_{j\in\mathcal{I}_{i}}\sum_{k>j}W_{ij}W_{ik}Z_{j}(t)Z_{k}(t).$
+    We apply a nonlinear function $f(\cdot)$ to the linear influence, such as the softplus function:
+    $f(x)=\ln(1+e^{x}).$
+    The total influence on channel $i$ at time $t$ is then given by:
+    $S_{i}(t)=f\left(L_{i}(t)\right)+\omega Q_{i}(t),$
+    where $\omega$ is a scaling parameter controlling the magnitude of the nonlinear interaction influence.
+    We update the initial channels with the total influence:
+    $\tilde{Z}_{i}(t)=Z_{i}(t)+S_{i}(t).$
+    III-C Removing Influences
+    III-C1 Removing Influences Using sDPCA
+    To isolate the channels of interest from the influences of the connected nodes, we apply sDPCA to the time series data of the connected nodes $\{Z_{j}(t)\}_{j\in\mathcal{C}}$.
+We obtain principal component scores $\{Z_{\text{pc},k}(t)\}_{k=1}^{N_{\text{dpc}}}$, where $N_{\text{dpc}}$ is the number of dynamic principal components retained.
+For each adjusted initial channel $\tilde{Z}_{i}(t)$, we regress it on the principal component scores and their interaction terms:
+    $\tilde{Z}_{i}(t)=\sum_{k=1}^{N_{\text{dpc}}}\varphi_{ik}Z_{\text{pc},k}(t)+\sum_{k\neq l}\psi_{ikl}Z_{\text{pc},k}(t)Z_{\text{pc},l}(t)+\epsilon_{i}(t).$
+    We then use the residuals $\hat{\epsilon}_{i}(t)$ as the adjusted time series for further analysis:
+    $\hat{Z}_{i}(t)=\hat{\epsilon}_{i}(t).$
+    III-C2 Granger Causality Testing
+    Using the adjusted time series $\{\hat{Z}_{i}(t)\}_{i=1}^{N}$, we fit vector autoregressive (VAR) models and perform Granger causality tests to identify causal relationships among the initial channels.
+The effectiveness of the sDPCA adjustment is evaluated by comparing the detected causal relationships with the ground truth.
+    III-D Simulation Results
+    To evaluate the performance of the proposed method, we compute accuracy metrics based on the detection of true causal relationships.
+Specifically, we compare the identified causal links with the known ground truth and compute measures such as accuracy rate.
+    By conducting both linear and nonlinear simulations, we assess the robustness of the sDPCA combined with Granger causality testing in high-dimensional settings with complex interdependencies.
+The results provide insights into the method’s effectiveness in isolating the channels of interest and accurately detecting causal relationships, even in the presence of nonlinear influences.
     The results are sensitive to the choice of latent processes and the influence of other channels.
 Different scenarios are explored by varying the number of influencing channels and the weight of their influence, with the goal of assessing the robustness and reliability of the proposed method in detecting true causal relationships within high-dimensional time series networks.
-For example, Table 1 presents the accuracy results for one of the simulation studies for different numbers of influencing nodes and varying weights on the COI in a high-dimensional network.
-The COI consists of the first 20 nodes out of a total of 256 nodes.
-Influencing Nodes = 10
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.90 0.85 0.60 0.35
-6 1.00 0.90 0.65 0.40
-12 0.85 0.85 0.60 0.45
-24 0.85 0.75 0.70 0.25
+For example, Table I presents the accuracy results for one of the simulation studies for different numbers of influencing nodes and varying weights on the COI in a high-dimensional network.
+The COI consists of the first 40 nodes out of a total of 512 nodes.
+The results of the other 5 simulation studies are given in the appendix.
+    TABLE I: Accuracy results for varying numbers of influencing nodes with nonlinear influence and weights on the COI in a high-dimensional network under a non-normal error distribution, $\epsilon\sim\mathrm{non\text{-}Norm}(\mu=-1,\sigma=1,\gamma=1.5,\kappa=3.5).$ The COI consists of the first 40 nodes out of a total of 512 nodes.
+The numbers 30, 60, 90, and 120 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, and 0.7 represent the weights of this influence.
     Influencing Nodes = 30
     Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.90 0.35 0.15 0.30
-6 0.85 0.80 0.30 0.25
-12 1.00 0.85 0.40 0.40
-24 0.95 0.60 0.35 0.20
+    3 0.96 0.88 0.61 0.65
+    6 0.97 0.77 0.62 0.64
+    12 0.97 0.82 0.61 0.65
+    24 0.93 0.82 0.65 0.64
     Influencing Nodes = 60
     Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.80 0.55 0.25 0.20
-6 0.70 0.35 0.20 0.20
-12 0.95 0.45 0.20 0.45
-24 0.90 0.10 0.30 0.20
+    3 0.98 0.66 0.64 0.64
+    6 0.99 0.73 0.64 0.65
+    12 0.94 0.69 0.64 0.64
+    24 0.96 0.67 0.64 0.64
     Influencing Nodes = 90
     Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.90 0.30 0.20 0.15
-6 0.85 0.15 0.30 0.20
-12 0.80 0.25 0.30 0.30
-24 0.85 0.40 0.35 0.25
-Table 1: Accuracy results for varying numbers of influencing nodes and weights on the COI in a high-dimensional network under a different non-normal error mechanism with different skewness and kurtosis parameters, $\epsilon\sim\mathrm{non-Norm}(\mu=0,\sigma=1,\gamma=0.5,\kappa=0.5).$ The COI consists of the first 20 nodes out of a total of 256 nodes.
-The numbers 10, 30, 60, 90 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, 0.7 represent the weights of this influence.
-This simulation evaluates the performance of the sDPCA method with non-normal error terms, adding complexity due to potential spurious causal relationships.
-The study employs permutation-based Granger causality testing, detailed in Section 2, to robustly evaluate causal relationships without assuming normally distributed errors.
-With 10 influencing nodes, sDPCA shows high accuracy, especially at lower weights (0.1), achieving 0.90 accuracy with 3 Ndpc.
-However, accuracy declines with higher weights (0.5 and 0.7), dropping to 0.60 and 0.35, respectively, indicating challenges in isolating true causal associations under heavier influences.
-As the number of influencing nodes increases to 30, a similar pattern emerges.
-High accuracy persists at lower weights (0.1), but significantly decreases with higher weights, dropping to 0.30 at a weight of 0.7, underscoring the difficulty in maintaining accuracy with greater external influence.
-With 60 and 90 influencing nodes, the challenge is more pronounced.
-For instance, with 60 influencing nodes and a weight of 0.7, accuracy falls to 0.20, highlighting the significant impact of increased external influences on the COI, complicating the isolation and accurate determination of true causal relationships.
-Overall, this table shows that while the sDPCA method can effectively isolate the COI and maintain high accuracy in the presence of external influences, its performance is sensitive to the weight of the influence and the number of influencing nodes.
-Lower weights and fewer influencing nodes generally result in higher accuracy, while higher weights and more influencing nodes pose greater challenges.
-These findings are consistent with the theoretical expectations discussed in our simulation framework, where the ability to isolate COI and accurately assess causal interactions is influenced by the degree of external interference.
-The use of permutation-based GC testing ensures that the assessment of causality is robust and not biased by the assumptions of normality, providing a more reliable analysis of the true causal relationships in the network.
-Tables for additional simulations using different error mechanisms to generate the time series are provided in the appendix.
-4 Application to EEG data
-In this section, we demonstrate the utility of the proposed approach.
-For practical purposes, we focus on one epoch of the data and analyze four channels: AF3, AF4, C3, and C4.
-The dataset comes from an experiment involving a series of motor/imagery tasks, providing rich insights into the temporal dynamics of brain activity.
-The original EEG recordings were collected from 109 volunteers using the BCI2000 system with a 64-channel setup, and the data is publicly available at PhysioNet.org (Schalk et al., 2004).
-The experiment included 14 runs for each subject: two baseline runs (one with eyes open and one with eyes closed) and twelve task-oriented runs.
-During the task runs, participants either performed or imagined movements in response to visual cues, such as opening and closing fists or feet.
-The tasks were divided into categories: physical movement of the left or right fist, imagined movement of the left or right fist, physical movement of both fists or both feet, and imagined movement of both fists or both feet.
-Our main objective is to observe the effect of the sDPCA method on dimensionality reduction and its impact on the isolated channels.
-Specifically, we aim to determine whether this method distorts the temporal information in the time series.
-Consequently, we focus on only a few channels and share the results as an example, rather than conducting a detailed multichannel analysis.
-This allows us to provide a demonstration of the approach’s effectiveness in isolating channels of interest and assessing their causal interactions.
-In Figure 1, the obtained three principal component (PC) scores are plotted over the time series data from which they were derived.
-This illustrates that the PC scores can capture the temporal structures of the high-dimensional time series from which they originate.
-This indicates that the obtained scores preserve the main characteristics of the time series, accurately representing their dynamics.
-This is a necessary finding that emphasizes the effectiveness and reliability of the procedure.
+    3 0.96 0.63 0.63 0.65
+    6 0.94 0.62 0.64 0.64
+    12 0.95 0.62 0.64 0.64
+    24 0.91 0.64 0.64 0.64
+    Influencing Nodes = 120
+    Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
+    3 0.95 0.65 0.63 0.64
+    6 0.95 0.65 0.64 0.65
+    12 0.90 0.64 0.65 0.64
+    24 0.97 0.64 0.64 0.64
+    This simulation evaluates the performance of the sDPCA method under a nonlinear influence scheme with non-normal error terms, introducing complexity due to both nonlinearity and deviations from normality.
+    At a low influence weight (weight = 0.1), sDPCA maintains high accuracy across all numbers of influencing nodes, even as the network becomes super connected and super dense.
+Accuracy exceeds 0.90 when the number of influencing nodes increases from 30 up to 120, indicating strong robustness in isolating the channels of interest (COI) even in highly interconnected networks where each node affects the causal pairs.
+    As the influence weight increases to 0.3, accuracy decreases but remains relatively high with fewer influencing nodes.
+For example, with 30 influencing nodes and Ndpc = 3, the accuracy is 0.88.
+However, as the network becomes denser (influencing nodes ¿ 30), the decline in accuracy becomes more pronounced, reflecting the increased challenge of isolating true causal relationships in a super connected network.
+    At higher influence weights (weight = 0.5 and 0.7), accuracy declines further and stabilizes around 0.64–0.65, regardless of the number of influencing nodes or the number of dynamic principal components (Ndpc).
+This plateau suggests a limit to the method’s ability to isolate the COI under strong nonlinear influences and pervasive connectivity, as heavier weights and a denser network make it more challenging to distinguish true causal relationships from spurious ones introduced by nonlinearity and widespread influence.
+    The impact of increasing Ndpc on accuracy is minimal at higher weights and in denser networks, indicating that adding more principal components does not enhance performance under these conditions.
+This suggests that the sDPCA method’s capacity to capture and remove complex influences is limited when the network is both highly connected and subject to strong nonlinear interactions.
+    Overall, the sDPCA method demonstrates robustness to nonlinear influences and non-normal error distributions at lower influence weights and in less densely connected networks.
+Its effectiveness diminishes as the strength of the nonlinear influence increases and as the network becomes super connected, highlighting the importance of considering both the magnitude of external influences and network connectivity when applying sDPCA in practical settings with nonlinear dynamics.
+    IV Application to EEG data
+    In this section, we apply our proposed approach to real EEG data from an experiment involving motor execution and imagery tasks, offering insights into the temporal dynamics of brain activity.
+EEG recordings were collected from 109 volunteers using the BCI2000 system with a 64-channel setup; the data is publicly available at PhysioNet.org [18].
+Each subject completed 12 task-oriented runs.
+Participants either physically performed or imagined movements in response to visual cues on a screen, such as opening and closing fists or feet depending on the target’s position.
+The tasks were divided into four categories: physical movements of the left or right fist, imaginary movements of the left or right fist, physical movements of both fists or both feet, and imaginary movements of both fists or both feet.
+EEG data were sampled at 160 Hz, along with an annotation channel indicating rest or the onset of motion (real or imagined) for specific movements.
+The EEG montage followed the international 10-10 system, omitting certain electrodes (Nz, F9, F10, FT9, FT10, A1, A2, TP9, TP10, P9, and P10).
+    All data underwent pre-processing to filter out artifacts like eye blinks and muscle movements, preserving the temporal structure of the recordings.
+The procedure included a high-pass filter at 0.5 Hz to eliminate low-frequency noise, removal of 60 Hz AC line noise, and Independent Component Analysis (ICA) using MNE-Python routines [19, 20] to isolate and exclude artifact components.
+Finally, artifacts associated with eye blinking and muscle movements were identified and removed, ensuring the recordings were free from these specific interferences.
+    Despite standard pre-processing, some recordings had unresolved issues like sawtooth traces, excessive blinking, broken channels, and unbalanced epochs.
+Therefore, we limited our analysis to 35 right-handed subjects (17 females and 18 males) whose recordings were qualified.
+To demonstrate how the proposed method can dissect causal connections between pairwise nodes in any time series network, we focused on six channels of interest (COI): AF3 and AF4 positioned over the pre-frontal cortex, C3 and C4 placed over the motor cortex, and O1 and O2 over the occipital cortex, as illustrated in Figure 1.
+These channels were chosen because they correspond to brain regions likely to be active during voluntary movement, imagery, and visual perception.
+    On the other hand, another important objective is to monitor the effect of the dimension reduction utilizing the sDPCA and its impact on the isolated channels.
+In particular, we propose that basic statistical tools such as trace plots, serial and partial autocorrelation functions, spectrum estimates, and many more can be practically used to determine the extent to which the method alters time-dependent properties in time series.
+This helps us to assess the efficacy of the proposed methodology for isolating COI while maintaining the temporal characteristics of the signals.
+    In Figure 1, we illustrate an example of one epoch from the dataset.
+The obtained three principal component (PC) scores are plotted over the time series data from which they were derived.
+This illustrates that the PC scores can capture the temporal structures/dynamics of the high-dimensional time series from which they originate and preserve the main characteristics of the time series.
     Figure 1: Isolating Nodes via Spectral Dynamic PCA method
-Within the scope of this study, we compared the conditions before and after applying the spectral dynamic principal component analysis (sDPCA) method to the channels of interest.
-As illustrated in Figures 2 to 3, although the temporal characteristics of the series remain unchanged, there is a noticeable trend toward stationarity in the serşal correlation amplitudes, which initially showed signs of non-stationarity.
+    We compared the time-dependent statistical properties before and after applying the spectral dynamic principal component analysis (PCA) method to the channels of interest.
+As illustrated in Figures 2, 3 and 4, although the temporal characteristics of the series maintained, there is a noticeable trend toward stationarity in the serial correlation amplitudes, which showed signs of non-stationarity before.
 This adjustment indicates that while the sequential properties of the series are preserved, they become more aligned with the assumptions required by the auxiliary model used for Granger causality (GC) testing.
+    Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+    (a) trace (b) acf (c) pacf (d) spectrum
     Figure 2: Sample vs Isolated channels C3 and C4.
+    Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+    (a) trace (b) acf (c) pacf (d) spectrum
     Figure 3: Sample vs Isolated channels AF3 and AF4.
-Moreover, our test results, summarized in Table 1, reveal significant differences in GC tests conducted before and after the isolation process.
-In several instances, the GC tests showed contrasting results after the isolation, underlining the presence of spurious GC causality in the original analysis.
-Specifically, we observed that causal relationships detected in the pre-isolation data were not present in the post-isolation data.
-This suggests that the influences of other channels, when not accounted for, can lead to misleading conclusions about causality.
-Direction of GC GC result Before GC result After
-C3 $\rightarrow$ AF3 Cause Do not cause
-C3 $\rightarrow$ AF4 Cause Cause
-C4 $\rightarrow$ AF3 Do not cause Do not cause
-C4 $\rightarrow$ AF4 Cause Cause
-AF3 $\rightarrow$ C3 Cause Do not cause
-AF3 $\rightarrow$ C4 Cause Cause
-AF4 $\rightarrow$ C3 Cause Cause
-AF4 $\rightarrow$ C4 Do not cause cause
-Table 2: Comparison of Granger Causality Before and After Applying sDPCA
-Overall, these findings highlight the effectiveness of the sDPCA method in refining the data by removing extraneous influences, thereby providing a more accurate assessment of the true causal relationships within the EEG channels.
-This approach enhances the reliability of GC tests, ensuring that the detected causality is genuine and not a result of spurious interactions.
-The findings accentuate the capability of the sDPCA method to reduce the dimensionality of the data without losing temporal information.
-5 Discussion and Conclusion
-The study aims at improving the accuracy of Granger causal analysis in high-dimensional time-series networks, specifically focusing on EEG datasets.
-Through the use of the sDPCA method, we have been able to effectively isolate the COI from the influence of other channels within the network.
-The results of the simulation studies and the analysis of the real EEG signals show the simplicity, usefulness, and effectiveness of the proposed method.
-The simulation study illustrates the challenges of identifying true causal interactions due to potential spurious causality, especially given the presence of non-normality.
-Integrating sDPCA and permutation-based GC testing significantly improved the accuracy of causal inference.
-The results show that the method maintained high accuracy in the presence of a moderate number of influencing nodes.
-However, the method’s accuracy decreases as the number of influencing nodes increases or approaches super/over connectivity (i.e., swarm-like neuronal connectivity).
-These findings underscore the sensitivity of GC to external influences and stress the necessity of accurately modeling such complexities in real-world data.
-The application of the proposed approach to a network of EEG signals provided empirical validation of its potential utility.
+    Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+    (a) trace (b) acf (c) pacf (d) spectrum
+    Figure 4: Sample vs Isolated channels O1 and O2.
+    Test results given in Table II, reveal significant differences in GC tests conducted before and after the isolation process.
+In several instances, the GC tests showed contrasting results after the isolation (i.e., dimension reduction), underlining the presence of spurious GC regarding 64-dimensional VAR-based GC analysis.
+We observed that some of the causal relationships detected in the pre-isolation (i.e., 64-dimensional VAR) were not present in the post-isolation (i.e., 6-dimensional VAR).
+This suggests that the influences of other channels, when not addressed properly, can lead to misleading conclusions about causality.
+On the other hand, estimating causal relationships in networks with many time series —using VAR models and Granger Causality tests- is quite challenging, both statistically and computationally.
+The main issue is the ”curse of dimensionality.” As the number of variables increases, the number of parameters we need to estimate grows rapidly, leading to a paralyzing parameter space.
+This can result in overfitting, higher variance in estimators, and problems like multicollinearity when variables are highly correlated.
+Conventional estimation methods like Ordinary Least Squares (OLS) become unfeasible in this context.
+Additionally, selecting the right lag order and variables is more complicated with so many variables, increasing the risk of misspecification and inflated type I errors due to multiple testing.
+On the computational side, high-dimensional VAR models require significant resources that burst computational time when dealing with multiple epochs and subjects.
+This makes the Granger Causality testing process less reliable and efficient in these high-dimensional settings.
+    TABLE II: Comparison of GC Before and After Applying sDPCA
+    Direction of GC GC test w.r.t 64-dim VAR (Before) GC test w.r.t 6-dim VAR (After)
+    1C3 $\rightarrow$ C4 $\blacksquare$ $\blacksquare$
+    1C3 $\rightarrow$ AF3 $\blacksquare$ $\square$
+    1C3 $\rightarrow$ AF4 $\square$ $\blacksquare$
+    1C3 $\rightarrow$ O1 $\blacksquare$ $\blacksquare$
+    1C3 $\rightarrow$ O2 $\blacksquare$ $\square$
+    1C4 $\rightarrow$ C3 $\square$ $\blacksquare$
+    1C4 $\rightarrow$ AF3 $\square$ $\square$
+    1C4 $\rightarrow$ AF4 $\square$ $\blacksquare$
+    1C4 $\rightarrow$ O1 $\square$ $\square$
+    1C4 $\rightarrow$ O2 $\square$ $\blacksquare$
+    AF3 $\rightarrow$ AF4 $\square$ $\blacksquare$
+    AF3 $\rightarrow$ C3 $\square$ $\square$
+    AF3 $\rightarrow$ C4 $\square$ $\square$
+    AF3 $\rightarrow$ O1 $\square$ $\square$
+    AF3 $\rightarrow$ O2 $\square$ $\blacksquare$
+    AF4 $\rightarrow$ AF3 $\square$ $\blacksquare$
+    AF4 $\rightarrow$ C3 $\square$ $\square$
+    AF4 $\rightarrow$ C4 $\blacksquare$ $\blacksquare$
+    AF4 $\rightarrow$ O1 $\square$ $\square$
+    AF4 $\rightarrow$ O2 $\square$ $\square$
+    FO1 $\rightarrow$ O2 $\square$ $\blacksquare$
+    FO1 $\rightarrow$ AF3 $\blacksquare$ $\square$
+    FO1 $\rightarrow$ AF4 $\square$ $\square$
+    FO1 $\rightarrow$ C3 $\blacksquare$ $\blacksquare$
+    FO1 $\rightarrow$ C4 $\blacksquare$ $\square$
+    FO2 $\rightarrow$ O1 $\square$ $\blacksquare$
+    FO2 $\rightarrow$ AF3 $\square$ $\blacksquare$
+    FO2 $\rightarrow$ AF4 $\square$ $\blacksquare$
+    FO2 $\rightarrow$ C3 $\square$ $\blacksquare$
+    FO2 $\rightarrow$ C4 $\square$ $\square$
+    **Black squares, $\blacksquare$, indicate a ”Cause”, while white squares, $\square$, represent a ”Do not Cause” relationship.
+**Black squares, $\blacksquare$, indicate a ”Cause”, while white squares, $\square$, represent a ”Do not Cause” relationship.
+    In this section, we also estimate and present the connectivities for COI in the context of data from 35 subjects pertaining to both motor execution (ME) and motor imagery (MI) tasks.
+The subjects performed a series of opening and closing (OC) gestures, including left-hand fist (LH), right-hand fist (RH), both fists’ (BH), and both feet’ (BF), for both the ME and MI tasks.
+The aim was to compare the brain activity associated with these tasks and gestures by identifying the patterns of connectivity and causal interactions amongst the selected channels.
+Each subject repeated the relevant task (i.e., task1: OC of LH/RH, task2: imagery OC of LH/RH, task3: OC of BH/BF, task4: imagery OC of BH/BF) 3 times (i.e., runs).
+Within these three task-oriented runs, they performed each motor movement/imagery motion (i.e., LH, RH, BH, BF) approximately 23 times in total.
+    Each 64 dimensional epoch consists of about 4 seconds, that is, 640 observations.
+Initially, the relevant epochs were extracted from the dataset with the help of annotation information.
+In order to simulate a high-dimensional space for 64-dimensional EEG, we augmented the real dataset to include 170 additional signals to 64 channels by creating additional features.
+By increasing the dimension of each epoch, we also aimed to enrich the dataset to better capture the underlying neural dynamics.
+This was achieved through a series of data augmentation steps designed to capture various aspects of neural activity and connectivity.
+First, we generated hemispheric asymmetry channels by calculating the difference between pairs of electrodes located symmetrically across the left and right hemispheres (e.g., Fp1 - Fp2).
+Next, we computed anterior-posterior gradient channels by taking the difference between frontal and posterior electrode pairs (e.g., Fz - Pz).
+We also created lateral connectivity channels by finding differences between electrodes along the sides of the brain (e.g., FT7 - TP7).
+To examine central brain activities, we calculated midline activity channels by subtracting posterior midline electrodes from anterior ones (e.g., Fpz - Oz).
+Additionally, we computed regional average channels by averaging the signals from electrodes within specific brain regions—frontal, central, parietal, and occipital.
+We employed spatial derivative channels (Laplacian) by subtracting the average signal of neighboring electrodes from the signal at a central electrode (e.g., Cz minus the average of FCz, CPz, C1, and C2).
+This enhances the detection of localized neural activity by emphasizing differences from the surrounding area.
+Furthermore, we introduced product interactions between specific pairs of electrodes by multiplying their signals (e.g., F3 times P3).
+This captures interactions (i.e., possibly nonlinear) and joint activations between different brain regions, which might be significant for complex cognitive processes.
+Then, for each epoch, the high-dimensional data were reduced to COI dimensions by the proposed approach, and the estimated connectivities were recorded.
+From these results, the connectivities shared by at least 70% of the subjects were determined and demonstrated in Figure 5.
+    (a) Connectivity at Left Fist MI
+    (b) Connectivity at Left Fist ME
+    (c) Connectivity at Right Fist MI
+    (d) Connectivity at Right Fist ME
+    (e) Connectivity at Both Fists MI
+    (f) Connectivity at Both Fists ME
+    (g) Connectivity at Both Feet MI
+    (h) Connectivity at Both Feet ME
+    Figure 5: Connectivity at MI vs Connectivity at ME.
+    The GC results revealed distinct patterns of connectivity among the six channels of interest during different motor tasks, as illustrated in Figure 5.
+When examining left-hand motor imagery, we observed that channel C3 (left motor cortex) G-causes channel C4 (right motor cortex) and channel O1 (left occipital cortex).
+Channel C4 also G-causes C3, AF4 (right pre-frontal cortex), O1, and O2 (right occipital cortex).
+This suggests that during the imagination of left-hand movement, there is significant bilateral interaction between motor cortices and connections to occipital regions involved in visual processing.
+    In contrast, during left-hand motor execution, the connectivity increased.
+Channel C3 G-causes C4, O1, and O2, indicating that ME engages both occipital cortices more than imagery does.
+The involvement of both visual regions during execution may reflect the integration of sensory feedback when performing physical movements.
+Similarly, for right-hand tasks, during motor imagery, channel C3 G-causes C4, O1, and O2, and channel C4 G-causes C3, AF3 (left pre-frontal cortex), O1, and O2.
+This pattern highlights strong bilateral connectivity and suggests that both hemispheres are actively engaged even during imagined movements.
+    When comparing motor imagery and execution for both fists, the connectivity patterns showed that channel C4 had extensive Granger causality to other channels during imagery, indicating the right motor cortex’s central role in coordinating bilateral imagined movements.
+During execution, there was increased causality from AF3 and AF4 (pre-frontal cortices) to motor and occipital regions, suggesting greater involvement of executive functions and planning during actual movement.
+For both feet tasks, the connectivity was less pronounced, with fewer significant Granger causalities observed during both imagery and execution.
+This may imply that foot movements engage different neural pathways or that the linear assessment is insufficient to capture the complexity.
+    Overall, the contrasts between motor imagery and execution reveal that actual movement involves more extensive and complex connectivity, particularly with increased involvement of pre-frontal regions during execution.
+    In conclusion, the proposed approach is an easy-to-implement method for examining the connectivity between a few nodes in a high-dimensional setting.
+By efficiently isolating the nodes of interest and removing the confounding effects of other channels, one may focus on specific neural interactions without the computational burden of analyzing the entire high-dimensional network.
+    V Discussion and Conclusion
+    Our study aimed to enhance the accuracy of Granger causality analysis in high-dimensional time-series networks, focusing on EEG datasets.
+By using the sDPCA method, we effectively isolated the channels of interest from the influence of other channels in the network.
+Results from both simulation studies and real EEG signal analysis demonstrate the simplicity and effectiveness of our proposed method.
+    We found that the method maintained high accuracy with a moderate number of influencing nodes.
+However, its accuracy decreased as the number of influencing nodes increased or when the network approached super connectivity, like swarm-like neuronal connections.
+This highlights the sensitivity of Granger causality to external influences and emphasizes the need to accurately model such complexities in real-world data.
+    Applying the proposed approach to EEG signal networks provided empirical validation of its utility.
 The principal component scores derived from the spectral domain accurately captured the temporal dynamics of the high-dimensional signals.
-The results indicate a convergence towards stationarity in the serial correlation amplitudes after applying sDPCA, getting the data closer to the assumptions demanded for reliable GC testing.
-The comparison of pre- and post-isolation GC results reveals the presence of spurious causality identification in the pre-isolation data, underlining the need to remove extraneous influences for accurate causality assessment.
-Despite promising findings, there are several limitations to our study.
-Although comprehensive, the simulation framework may only capture some nuances of real EEG data.
-In addition, the test results can be significantly affected by the choice of hyper-parameters, such as the number of principal components, the weights, and the number of influencing nodes.
-These parameters need to be selected according to the specific characteristics of the datasets to be studied.
-Furthermore, our method’s reliance on linearity-based PCA in the frequency domain assumes that the underlying temporal dynamics can be adequately captured, which may not always be the case.
-Future research should consider the integration of sDPCA with the perspective of capturing non-linear relationships.
-To further improve the robustness and accuracy of causal inference in complex networks, other dimensionality reduction techniques and advanced GC methods should be targeted.
-In addition, the applicability of the methodology could be significantly enhanced by extending it to include more sophisticated models of external influences and by incorporating real-time analysis capabilities.
-In conclusion, our proposed approach provides a practical and reliable solution for isolating and analyzing causal associations in high-dimensional time series networks.
-The ability of the method to accurately identify true causal connections while mitigating the effects of confounding influences makes it a valuable tool for researchers in neuroscience and other fields dealing with complex time series data.
-The promising results of the study suggest that this approach can significantly improve the reliability of causality analysis, paving the way for more accurate and insightful investigations of the causal dynamics of high-dimensional networks.
-Appendix A Appendix / supplemental material
-Influencing Nodes = 10
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.95 0.40 0.50 0.35
-6 0.95 0.40 0.35 0.30
-12 0.90 0.75 0.50 0.70
-24 0.90 0.65 0.25 0.75
-Influencing Nodes = 30
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.90 0.40 0.40 0.25
-6 0.90 0.40 0.25 0.40
-12 0.80 0.50 0.45 0.45
-24 0.80 0.35 0.45 0.40
-Influencing Nodes = 60
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.35 0.25 0.45 0.15
-6 0.85 0.30 0.50 0.20
-12 0.65 0.15 0.05 0.35
-24 0.40 0.50 0.25 0.15
-Influencing Nodes = 90
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.80 0.35 0.30 0.45
-6 0.55 0.30 0.25 0.15
-12 0.40 0.40 0.20 0.30
-24 0.30 0.30 0.20 0.20
-Table 3: Accuracy results for varying numbers of influencing nodes and weights on the COI in a high-dimensional network under a non-normal error mechanism with different causal structure, $\epsilon\sim\mathrm{non-Norm}(\mu=0,\sigma=1,\gamma=0.5,\kappa=0.5).$ The COI consists of the first 20 nodes out of a total of 256 nodes.
-The numbers 10, 30, 60, 90 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, 0.7 represent the weights of this influence.
-Influencing Nodes = 10
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.95 0.75 0.75 0.45
-6 0.85 0.75 0.45 0.05
-12 0.95 0.75 0.75 0.60
-24 1.00 0.90 0.60 0.50
-Influencing Nodes = 30
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.95 0.65 0.20 0.35
-6 0.90 0.35 0.50 0.20
-12 0.95 0.50 0.30 0.30
-24 1.00 0.60 0.15 0.40
-Influencing Nodes = 60
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.90 0.20 0.25 0.20
-6 0.85 0.50 0.20 0.40
-12 0.85 0.40 0.25 0.30
-24 0.90 0.15 0.15 0.35
-Influencing Nodes = 90
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.75 0.35 0.40 0.30
-6 0.90 0.25 0.30 0.15
-12 0.95 0.60 0.40 0.20
-24 0.90 0.25 0.35 0.10
-Table 4: Accuracy results for varying numbers of influencing nodes and weights on the COI in a high-dimensional network under non-normal error-generating mechanisms with different skewness and kurtosis parameters, $\epsilon\sim\mathrm{non-Norm}(\mu=-1,\sigma=1,\gamma=1.5,\kappa=3.5).$.
-The COI consists of the first 20 nodes out of a total of 256 nodes.
-The numbers 10, 30, 60, 90 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, 0.7 represent the weights of this influence.
-Influencing Nodes = 10
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.95 0.90 0.60 0.45
-6 0.90 0.90 0.50 0.45
-12 0.90 0.90 0.50 0.20
-24 1.00 0.95 0.70 0.55
-Influencing Nodes = 30
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.95 0.50 0.30 0.30
-6 0.85 0.55 0.30 0.20
-12 0.95 0.50 0.20 0.10
-24 0.85 0.65 0.45 0.25
-Influencing Nodes = 60
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.85 0.35 0.05 0.20
-6 0.75 0.30 0.25 0.35
-12 0.90 0.35 0.15 0.30
-24 0.80 0.55 0.45 0.25
-Influencing Nodes = 90
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.90 0.25 0.15 0.35
-6 0.85 0.40 0.20 0.25
-12 0.90 0.25 0.05 0.05
-24 1.00 0.20 0.20 0.25
-Table 5: Accuracy results for varying numbers of influencing nodes and weights on the COI in a high-dimensional network with Gaussian error-generating mechanisms.
-The COI consists of the first 20 nodes out of a total of 256 nodes.
-The numbers 10, 30, 60, 90 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, 0.7 represent the weights of this influence.
-Influencing Nodes = 10
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.95 0.65 0.25 0.35
-6 0.90 0.65 0.50 0.45
-12 0.95 0.55 0.55 0.25
-24 0.90 0.65 0.25 0.30
-Influencing Nodes = 30
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.90 0.55 0.20 0.10
-6 0.70 0.65 0.30 0.30
-12 0.75 0.55 0.50 0.35
-24 0.75 0.65 0.40 0.30
-Influencing Nodes = 60
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.80 0.50 0.35 0.15
-6 0.70 0.40 0.20 0.15
-12 0.90 0.45 0.20 0.30
-24 0.75 0.55 0.20 0.25
-Influencing Nodes = 90
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.55 0.25 0.35 0.30
-6 0.55 0.60 0.15 0.25
-12 0.65 0.20 0.30 0.25
-24 0.45 0.30 0.15 0.20
-Table 6: Accuracy results for varying numbers of influencing nodes and weights on the COI in a high-dimensional network with another Gaussian error-generating mechanism.
-The COI consists of the first 20 nodes out of a total of 256 nodes.
-The numbers 10, 30, 60, 90 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, 0.7 represent the weights of this influence.
-Influencing Nodes = 10
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 1.00 0.60 0.70 0.30
-6 0.90 0.40 0.40 0.35
-12 0.95 0.15 0.25 0.30
-24 1.00 0.40 0.45 0.70
-Influencing Nodes = 30
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.85 0.35 0.30 0.30
-6 0.70 0.40 0.20 0.40
-12 0.75 0.35 0.40 0.25
-24 0.80 0.50 0.50 0.40
-Influencing Nodes = 60
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.40 0.50 0.35 0.15
-6 0.50 0.70 0.30 0.25
-12 0.50 0.40 0.25 0.25
-24 0.40 0.60 0.20 0.40
-Influencing Nodes = 90
-Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-3 0.45 0.25 0.20 0.25
-6 0.45 0.25 0.15 0.40
-12 0.35 0.55 0.25 0.30
-24 0.60 0.30 0.40 0.55
-Table 7: Accuracy results for varying numbers of influencing nodes and weights on the COI in a high-dimensional network with another Gaussian error-generating mechanism.
-The COI consists of the first 20 nodes out of a total of 256 nodes.
-The numbers 10, 30, 60, 90 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, 0.7 represent the weights of this influence.
+After applying sDPCA, we observed a convergence towards stationarity in the serial correlation amplitudes, bringing the data closer to the assumptions required for reliable Granger causality testing.
+Comparing the results before and after isolation revealed spurious causality identifications in the pre-isolation data, underlining the importance of removing extraneous influences for accurate causality assessment.
+    Despite these promising findings, there are some limitations to our study.
+The simulation framework, while comprehensive, may not capture all nuances of real EEG data.
+Additionally, the test results can be significantly affected by the choice of hyperparameters, such as the number of principal components, the weights, and the number of influencing nodes.
+These parameters need to be carefully selected based on the specific characteristics of the datasets under study.
+Furthermore, our method relies on linearity-based PCA in the frequency domain, assuming that the underlying temporal dynamics can be adequately captured, which may not always hold true.
+Future research should consider integrating sDPCA with approaches that capture nonlinear relationships.
+To further improve the robustness and accuracy of causal inference in complex networks, exploring other dimensionality reduction techniques and advanced Granger causality methods would be beneficial.
+Moreover, the applicability of the methodology could be enhanced by including more sophisticated models of external influences and by incorporating real-time analysis capabilities.
+    Overall, our proposed approach offers a practical solution for isolating and analyzing causal associations in high-dimensional time series networks.
+Its ability to accurately identify causal connections while mitigating the effects of confounding influences makes it a valuable tool for researchers in neuroscience and other fields dealing with complex signals.
+The encouraging results of the study suggest that this method can improve the reliability of causality analysis, leading to more accurate and insightful investigations of the causal dynamics in high-dimensional networks.
+    VI Acknowledgments
+    We acknowledge the assistance of Grammarly and DeepL AI tools for grammar checks and language enhancement during the preparation of this manuscript.

### v3

@@
 Abstract
-A novel approach is developed for discovering directed connectivity between specified pairs of nodes in a high-dimensional network (HDN) of brain signals.
-To accurately identify causal connectivity for such specified objectives, it is necessary to properly address the influence of all other nodes within the network.
-The proposed procedure herein starts with the estimation of a low-dimensional representation of the other nodes in the network utilizing (frequency-domain-based) spectral dynamic principal component analysis (sDPCA).
-The resulting scores can then be removed from the nodes of interest, thus eliminating the confounding effect of other nodes within the HDN.
+A novel and practical approach is introduced for identifying causal connectivity between specific node pairs in high-dimensional networks (HDN) of brain signals.
+The novelty lies in enabling a conventional Granger causality analysis to work effectively in a high-dimensional setting by combining it with frequency-domain-based spectral dynamic principal component analysis (sDPCA).
+The proposed procedure begins with the estimation of a low-dimensional representation of the other nodes in the network utilizing sDPCA.
+The resulting scores can then be extracted from the nodes of interest, thus removing the confounding effect of other nodes within the HDN.
 Accordingly, causal interactions can be dissected between nodes that are isolated from the effects of the network.
-Extensive simulations have demonstrated the effectiveness of this approach as a tool for causality analysis in complex time series networks.
-The proposed methodology has also been shown to be applicable to multichannel EEG networks.
-Index Terms:
-High-Dimensional Networks of Signals, Connectivity, Causality, Spectral Dynamic PCA, EEG signals
-I Introduction
-Identifying effective connectivity in high-dimensional networks (HDN) is a challenging problem that remains an active area of research and has been previously addressed by several studies [see, e.g., 1, 2, 3, 4, 5, 6, 7, 8].
-Moreover, it is difficult to infer or interpret interactions between a pair of nodes while taking into account potential interactions with all other nodes in the background [9, 10].
-This is a particularly relevant issue (e.g., analysis of EEG-like signals) when the focus might stand on a few channels/nodes of interest (COI/NOI) within an extensive network of interacting nodes.
-In fact, utilizing conventional vector autoregressive models (VAR) can be a practical way to identify pairwise causal interactions in high-dimensional networks as long as the potential confounding effect due to the high-dimensional background nodes is addressed.
-In this paper, we propose a framework that provides a low-dimensional summary of the background activity in the high-dimensional network (HDN).
-To discern effective connectivity, particularly between pairs of nodes in an HDN of brain signals, we propose a practical and easily accessible approach based on spectral dynamic principal component analysis (PCA) ([11, 12]) combined with testing for GC in lieu of the conventional F-test.
-The sDPCA method enhances the classical PCA by incorporating temporal dependency within the data, thereby reducing the dimensionality of time series in the frequency domain while retaining maximal temporal information.
-In other words, while the classical PCA is based on the eigendecomposition of the zero-lag (or contemporaneous) covariance matrix, sDPCA is based on the eigendecomposition of the spectral matrix at all frequencies.
-Since there is a 1-1 relationship between the spectral matrix (at all frequencies) and the covariance matrix (at all lags) then this approach captures lead-lag relationships amongst signals.
-Various methodologies have evolved from sDPCA.
-For instance, [13] developed a spectral envelope using sDPCA to identify common signals across multiple time series, while [14] proposed a functional variant of sDPCA for analyzing functional time-series data.
-As indicated by [15, 5, 16], spectral principal component analysis (sPCA) in the frequency domain offers practical benefits and is highly effective in low-dimensional representations of high-dimensional time series.
-The main objective of this study is to illustrate how the proposed approach enables the practical and sufficient isolation of COI by partialling out the influence of disinterested nodes, which are derived as a dynamic summary in a low-dimensional representation.
-The frequency-domain dynamic PCA method maintains the spectral characteristics of the data, which are crucial for accurately determining causal relationships in time series networks.
-The proposed approach offers two main advantages.
-Firstly, sDPCA in the frequency domain offers an accurate and practical dimension reduction while preserving the spectral information.
-This is essential because the spectrum of a signal captures cross-covariances at all lags.
-This leads to a more precise isolation of COI, reducing the likelihood of spurious causal inferences.
-Secondly, the proposed approach enables the formation of a comprehensive yet interpretable representation of the interactions between specified nodes at a defined level of resolution rather than being unable to infer by the expansive number of possible interactions between nodes.
-This facilitates the study of considerable numbers of interactions, ranging from pairs of nodes to interactions between a tractable number of clusters of nodes.
-The subsequent sections will elaborate on the proposed approach and its components.
-Section 2 presents the proposed approach.
-Section 3 outlines the scope of the simulation study and presents related results.
-Section 4 applies the proposed approach to EEG data and discusses the findings.
-II Methodology
-II-A Inference in high-dimensional networks of time series
-Define $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ to be a high dimensional network of signals with possibly $n\geq 256$ nodes.
+Simulations confirm that the approach reliably retrieves causal links, even in highly interactive networks.
+The method also performs well on multichannel EEG data, and findings align with earlier findings in the literature.
+Keywords: High-Dimensional Networks of Signals, Connectivity, Causality, Spectral Dynamic PCA, EEG signals
+1 Introduction
+Identifying effective connectivity in high-dimensional networks (HDN) remains a demanding problem and has been addressed by several studies (see, e.g., Kalisch and Bühlmann, 2014; Wang et al., 2016; Ting et al., 2016, 2017; Wang et al., 2019; Zarghami and Friston, 2020; Siggiridou and Kugiumtzis, 2021; Shojaie and Fox, 2022).
+Moreover, it is difficult to infer or interpret interactions between a pair of nodes while taking into account potential interactions with all other nodes in the background (Wang and Drton, 2020; Basu et al., 2024).
+One particular challenge arises when the objective is simply to identify the causal interaction between specifically interested pairs of nodes or pairs of subsets of nodes/channels within an extensive network of interacting nodes.
+Attempting to model an entire HDN for such a specific objective will require fitting a highly parameterized model with potentially hundreds of parameters, which can lead to over-parameterization of the task and interpretation difficulties.
+In fact, utilizing conventional bivariate vector autoregressive models (VAR) can be a practical and manageable way to identify pairwise causal interactions in high-dimensional networks as long as the potential confounding effect due to the high-dimensional background nodes is addressed.
+Adopting this strategy, we propose a procedure that uses conventional Granger causality (GC) testing to detect causal connections between any two nodes while accounting for the global background activity in a low-dimensional representation.
+Such a representation is constructed through spectral dynamic principal component analysis (sDPCA) (Brillinger, 1969, 1981), which extends the principal component analysis to incorporate temporal dependence.
+By working with the spectral density matrix —directly linked to all auto and cross-covariances- sDPCA preserves frequency-specific features critical for neural data analysis.
+The sDPCA enhances the PCA by incorporating temporal dependency within the data, thereby reducing the dimensionality of time series in the frequency domain while retaining temporal information.
+In other words, while the conventional PCA is based on the eigendecomposition at the zero-lag (or contemporaneous) covariance matrix, sDPCA is based on the eigendecomposition of the spectral density matrix estimated for lagged auto-covariances at all frequencies.
+Since there is a 1-1 relationship between the spectral matrix (at all frequencies) and the covariance matrix (at all lags), this approach captures lagged cross-correlations amongst signals.
+Various methodologies have been derived from frequency domain PCA.
+For instance, Stoffer et al.
+(1993); Stoffer (1999); Stoffer et al.
+(2000) developed a spectral envelope using dynamic PCA to identify common signals in multiple categorical and real-valued time series, Hörmann et al.
+(2015) proposed a variant of dynamic PCA for analyzing functional time-series data.
+As demonstrated by Ombao and Ho (2006); Wang et al.
+(2019); Ombao and Pinto (2022), principal component analysis in the frequency domain offers practical benefits and is highly effective in summarizing high-dimensional signals by capturing their main modes of variation.
+On the other hand, given the multivariate nature of the objectives outlined herein, it becomes relevant to inquire about the applicability of the state-of-the-art techniques of multivariate time series analysis.
+For example, Factor-augmented VAR (FAVAR) models (Bernanke et al., 2005; Lin and Michailidis, 2020) reduce dimensionality by extracting a handful of static factors from large panels and then placing these factors into a more manageable VAR system.
+This strategy centers on global modeling and treating the extracted factors as regressors to explain multiple target variables; however, it is not designated for cases where attention is on a single pair of nodes within a high-dimensional network.
+Dynamic factor models (DFMs) (Forni et al., 2000, 2015; Stock and Watson, 2005; Barigozzi and Hallin, 2017) similarly rely on a few latent components, focusing on decomposing each series into common and idiosyncratic components.
+DFM allows more detailed lagged structures than static factor approaches but often seeks to explain broad data variance or perform large-scale forecasts.
+Conditional Granger causality (CGC) (Geweke, 1984) instead tests whether adding one predictor refines forecasts of another, holding all others fixed.
+Although CGC reveals causal associations, it can become complicated in high dimensions when many variables must be jointly conditioned.
+While there are some parallels between the proposed approach and that of others, each method is characterized by a distinct emphasis.
+The proposed approach diverges by isolating specific node pairs.
+It accounts for background effects by spectral components and then applies GC testing on a simplified bivariate system.
+This diverges from CGC’s exhaustive conditioning on many channels, from FAVAR’s global factor plus VAR, and dynamic factor models’ emphasis on large-scale shared patterns.
+The approach also provides feasibility in large networks, where a smaller factor-driven representation is sufficient to identify causal interactions.
+The proposed approach distinguishes itself from existing approaches by focusing on frequency-domain dimensionality reduction combined with targeted pairwise causality analysis.
+The theoretical formulations of conditional GC, FAVAR, and DFMs have different aims – identifying influences through conditioning, incorporating large information sets into VARs, and explaining common variance in high dimensions, respectively – and each comes with certain assumptions about the data-generating process.
+The proposed approach does not compete with these state-of-the-art methods on their own terms; rather, it occupies a complementary cavity.
+It assumes that in many high-dimensional systems (e.g., brain recordings), one can extract the key oscillatory components and concentrate on their interactions, thereby enabling meaningful causal understanding without having to model every other interaction.
+Empirically, this translates to a convenient method that is sensitive to the features of brain monitoring data – handling many channels and frequency-specific processes – while building on the proven strengths of factor modeling and Granger causality.
+Given the objectives of established methods and our approach, it is important to note that utilizing GC with sDPCA does not propose a universal substitute for full network modeling.
+Instead, it is a domain-orientated alternative that emphasizes interpretability and feasibility in the analysis of high-dimensional network time series.
+This allows for targeted analysis, whether at the level of individual node pairs or among a tractable number of node clusters, without being constrained by finitely many interactions.
+In what follows, Section 2 presents the proposed approach for an HDN setting, Section 3 describes simulation setups and results, and Section 4 features an application to EEG recordings, showing how the method handles brain signals.
+2 Methodology
+2.1 Inference in high-dimensional networks of time series
+Define $V=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ to be a high dimensional network of signals with possibly $n\geq 256$ nodes.
 As we consider in this paper, this network may represent the activity of a brain through EEG recordings in which each component (i.e., node or channel) is actually a time series.
-In this notation of network $N$, $Y^{*}$ and $X^{*}$ denote any two nodes of interest.
-For problems whose solutions are sought component-wise, we set our goal as exploring the interaction or efficient connectivity between any pair of nodes in a network rather than the relationality of the whole network.
-These two nodes are subject to the influence of other nodes due to the network’s relational structure.
-This means that to dissect the respective efficient connection between any two nodes of interest; they need to be freed from the influences of other nodes.
+In this notation of network $V$, $Y^{*}$ and $X^{*}$ denote any two nodes of interest.
+For problems whose solutions are sought component-wise, we set our goal to explore the effective connectivity between any node pairs in a network rather than the associativity of the whole network.
+Node pairs are subject to the influence of other nodes due to the network’s relational structure.
+This means that to dissect the respective effective connection between any two nodes of interest; they need to be freed from the influences of other nodes.
 Thereby, the interference-free counterparts of the nodes of interest can be obtained in the following manner:
-$Y_{t}=Y^{*}_{t}-E\big{(}Y^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
-$X_{t}=X^{*}_{t}-E\big{(}X^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
+$\hat{Y}_{t}=Y^{*}_{t}-\hat{Y}^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t}),$
+$\hat{X}_{t}=X^{*}_{t}-\hat{X}^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t}),$
 where the $f_{z}(.)$ is a function that encompasses the dynamic variability of excluded nodes.
 Practically, the function $f_{z}(.)$ acts as a filter and can take any form as long as it provides the net summary discharge of the nodes excluded.
-Within the scope of this study, we utilize the spectral domain dynamic PCA implementation made available by Hörmann et al.
-(2015) via R package called freqdom.
-This method, originally proposed by Brillinger (1981), enables us to obtain principal component scores that concisely encapsulate variability in multi-dimensional time series, specifically with regard to their serial dependence [17, 16, see, e.g., ].
-Hence, we propose the filter $f_{z}(.)$ over excluded nodes to provide the following summary indicators,
-$f_{z}(Z_{1,t},\ldots,Z_{n-2,t})=\{pc1,pc2,\ldots,pc_{k},\text{interactions}\},$
-where $pc1,pc2,\ldots,pc_{k}$ are the scores of the principal components obtained through sDPCA, and interactions represent any interaction terms between these principal components.
-The conditional means with respect to these indicators are then given by:
-$E\big{(}Y^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=a_{1}\cdot pc1+a_{2}\cdot pc+$
-$\cdots+a_{k}\cdot pc_{k}+\sum a_{ij}\cdot(pc_{i}\times pc_{j}),$
-$E\big{(}X^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=b_{1}\cdot pc1+b_{2}\cdot pc2+$
-$\cdots+b_{k}\cdot pc_{k}+\sum b_{ij}\cdot(pc_{i}\times pc_{j}),$
-where $pc_{i}$ represents the score of the $i$-th dynamic principal component and $a_{i},b_{i},a_{ij},b_{ij}$ are coefficients corresponding to the principal components and their interactions.
-The number of principal components $k$ can be chosen to be relatively low compared to the number of dimensions, provided that they capture a sufficient proportion of the total variability in the network.
-For practical implementation, the first few principal components are generally sufficient.
+In the scope of this paper, the function $f_{z}(.)$ is intended to provide low-dimensional summaries of excluded nodes.
+To this end, we utilize the spectral-domain dynamic PCA implemented in the R package freqdom made available by Hörmann et al.
+(2015).
+Spectral PCA originally proposed by Brillinger (1981), enables us to obtain principal component scores that concisely encapsulate variability in multi-dimensional time series, specifically with regard to their serial dependence (Shumway et al., 2000; Ombao and Pinto, 2022, see, e.g., ).
+Hence, we propose the filter $f_{z}(\cdot)$ over the excluded nodes to yield a vector of summary indicators,
+$f_{z}(Z_{1,t},\ldots,Z_{n-2,t})=\mathbf{p}_{t}=\bigl{(}dpc_{1}(t),dpc_{2}(t),\ldots,dpc_{k}(t)\bigr{)},$
+where $dpc_{i}(t)$ denotes the score of the $i$-th dynamic principal component obtained through sDPCA.
+One may optionally augment $\mathbf{p}_{t}$ with higher-order transformations of the principal-component scores (for instance, cross-products $pc_{i}(t)\times pc_{j}(t)$ or polynomial expansions) if a more general functional form is required to capture nonlinearities or interactions among these components.
+In general, the conditional mean of $Y^{*}_{t}$ and $X^{*}_{t}$ given these indicators can be modeled as:
+$\hat{E}\bigl{(}Y^{*}_{t}\big{|}\mathbf{p}_{t}\bigr{)}=G_{Y^{*}}\bigl{(}\mathbf{p}_{t};\bm{\theta}_{Y^{*}}\bigr{)},$
+$\hat{E}\bigl{(}X^{*}_{t}\big{|}\mathbf{p}_{t}\bigr{)}=G_{X^{*}}\bigl{(}\mathbf{p}_{t};\bm{\theta}_{X^{*}}\bigr{)},$
+where $G_{Y^{*}}(\cdot;\bm{\theta}_{Y^{*}})$ and $G_{X^{*}}(\cdot;\bm{\theta}_{X^{*}})$ are functions (e.g.
+linear, nonlinear, or another parametric/nonparametric regression form) parameterized by $\bm{\theta}_{Y^{*}}$ and $\bm{\theta}_{X^{*}}$.
+Considering an option where the terms are linearly combined and have first-order interactions, the representation can be written as follows:
+$G_{Y^{*}}\bigl{(}\mathbf{p}_{t};\bm{\theta}_{Y^{*}}\bigr{)}=\alpha_{0}+\sum_{i=1}^{q}\alpha_{i}dpc_{i}(t)+\sum_{i=1}^{q}\sum_{j\neq i}^{q}\alpha_{ij}\bigl{[}dpc_{i}(t)\times dpc_{j}(t)\bigr{]},$
+Here, $q$ is typically small relative to the original dimension, so that $\{\alpha_{0},\alpha_{i},\alpha_{ij}\}$ remain tractable to estimate in practice.
 Spectral DPCA
@@
 […]
-The dynamic principal component scores $dpc_{t}^{(j)}$ are then calculated by filtering the time series $\mathbf{Z}_{t}$ with the dynamic principal component filters:
-$dpc_{t}^{(j)}=\sum_{k=-q}^{q}\phi_{k}^{(j)}\mathbf{Z}_{t-k}.$
-This convolution operation can be represented in matrix notation as:
-$\mathbf{dpc}=\mathbf{Z}\ast\mathbf{\Phi},$
-where $\ast$ denotes the convolution operation.
+By construction, sDPCA typically employs two-sided filters when forming the scores. This is seen in
+$dpc_{t}^{(j)}=\sum_{k=-q}^{q}\bigl{\{}\phi_{k}^{(j)}\bigr{\}}^{\!\top}\mathbf{Z}_{t-k},$
+where $\{\phi_{k}^{(j)}\}$ are the Fourier-based filter coefficients arising from the inverse transform of the $j$-th dynamic eigenvector.
+The index $k$ runs from $-q$ to $q$, reflecting a non-causal (two-sided) convolution that uses both past and future data relative to time $t$.
+This formulation can yield improved spectral estimates, though it assumes the entire data record is available.
+In real-time or causal applications, a one-sided variant can be adopted:
+$dpc_{t}^{(j)}=\sum_{k=0}^{q}\bigl{\{}\phi_{k}^{(j)}\bigr{\}}^{\!\top}\mathbf{Z}_{t-k},$
+which omits “future” values $\mathbf{Z}_{t+k}$ for $k>0$.
+This causal filter operates only on current and lagged observations, making it suitable for sequential or forecasting tasks.
+The choice between two-sided and one-sided implementations depends on whether retrospective or real-time usage is intended, as well as on the degree to which one wishes to exploit possible gains from including both past and future observations.
 The proportion of variance explained by the $j$-th dynamic principal component is quantified as:
@@
 These steps collectively facilitate the decomposition of multivariate time series into components that are uncorrelated in time and maximize the long-run variance explained.
-Terms and Notation
-  • $\mathbf{Z}$: Original multivariate time series matrix of dimensions $T\times n$, where $T$ is the number of time points and $n$ is the number of variables.
-  • $F_{\mathbf{Z}}(\omega)$: Cross-spectral density matrix at frequency $\omega$.
-  • $\text{Cov}(\mathbf{Z}_{h},\mathbf{Z}_{0})$: Covariance matrix at lag $h$.
-  • $\hat{C}_{\mathbf{Z}}(h)$: Empirical lagged auto-covariance matrix at lag $h$.
-  • $w(\cdot)$: Window function used to smooth the estimates.
-  • $q$: Window size for the kernel estimator.
-  • $\phi_{k}^{(j)}$: Dynamic principal component filters, obtained as the Fourier coefficients of the dynamic eigenvectors.
-  • $\varphi_{j}(\omega)$: Dynamic eigenvectors of the spectral density matrix $F(\omega)$.
-  • $dpc_{t}^{(j)}$: Dynamic principal component scores.
-  • $\mathbf{dpc}$: Matrix of dynamic principal component scores.
-  • $\mathbf{\Phi}$: Matrix of dynamic principal component filters.
-  • $\lambda_{j}(\omega)$: $j$-th dynamic eigenvalue of the spectral density matrix $F(\omega)$.
-II-B Granger Causality test
-Granger Causality (GC) is a statistical hypothesis test used to determine whether one time series can predict another.
+2.2 Granger Causality
+Granger Causality (GC) is a statistical concept used to determine whether one time series can predict another.
 The concept, introduced by Granger (1969), is based on the principle that if the prediction of a time series $Y_{t}$ can be improved by incorporating past values of another time series $X_{t}$, then $X_{t}$ is said to ”Granger-cause” $Y_{t}$.
-The Granger causality test involves the following steps. First, consider two time series $X_{t}$ and $Y_{t}$. To test if $X_{t}$ Granger-causes $Y_{t}$, we compare the following two models:
+The Granger causality analysis involves the following steps. First, consider two time series $X_{t}$ and $Y_{t}$. To test if $X_{t}$ Granger-causes $Y_{t}$, we compare the following two models:
 - Unrestricted Model:
@@
 This issue is particularly acute in complex systems with many interacting variables.
-II-C Proposed Approach
+2.3 Proposed Approach
 In high-dimensional networks of time series, analyzing causal connectivity between specific pairs of signals can be both challenging and computationally intensive due to the complexity of modeling interactions across all nodes.
@@
 Our approach utilizes spectral dynamic principal component analysis to capture the essential variability of the background network in a low-dimensional representation.
-By partialling out the influence of other nodes through sDPCA, we can isolate the COI and accurately assess the causal connectivity between them using conventional Granger causality tests.
+By partialling out the influence of other nodes through sDPCA, we can isolate the COI and accurately assess the causal connectivity between them using conventional Granger causality framework.
 This method simplifies the analysis by avoiding the need to model the entire high-dimensional network and enhances interpretability when the primary interest is in the causal relationship between specific signals.
 The proposed approach is given in the Procedure outline below.
-Input: High-dimensional Networks of Signals $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$
+Input: High-dimensional Networks of Signals $V=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$
 Output: Determination of causal relationship between $Y^{*}_{t}$ and $X^{*}_{t}$
-1. Identify Nodes of Interest: NOI:$\{Y^{*}_{t},X^{*}_{t}\}$.
-In practice, each node can also be a cluster of nodes that can be represented by sDPCA scores.
+1. Identify Nodes of Interest: NOI:$\{Y^{*}_{t},X^{*}_{t}\}$. In practice, each node can also be a cluster of nodes represented by sDPCA scores.
 2. Apply sDPCA:
   1. i
-    Estimate spectral density matrix $F_{\mathbf{Z}}(\omega)$ for $\mathbf{Z}_{t}=\{Z_{1,t},\ldots,Z_{n-2,t}\}$.
+    Estimate spectral density matrix $F_{\mathbf{Z}}(\omega)$ for $\mathbf{Z}_{t}$.
   2. ii
-    Compute dynamic eigenvectors $\varphi_{j}(\omega)$ and
-eigenvalues $\lambda_{j}(\omega)$.
-  • iii
-    Derive dynamic principal component filters:
-[…]
-  • iv
-    Calculate dynamic principal component scores:
-    $\text{dpc}_{t}^{(j)}=\sum_{k=-q}^{q}\phi_{k}^{(j)}\mathbf{Z}_{t-k}$
-  • 3. Partial Out Influence of Other Nodes:
-    $Y_{t}=Y^{*}_{t}-\hat{E}\big{(}Y^{*}_{t}\mid\text{dpc}_{t}^{(1)},\ldots,\text{dpc}_{t}^{(k)}\big{)}$
-    $X_{t}=X^{*}_{t}-\hat{E}\big{(}X^{*}_{t}\mid\text{dpc}_{t}^{(1)},\ldots,\text{dpc}_{t}^{(k)}\big{)}$
+    Compute dynamic eigenvectors $\varphi_{j}(\omega)$ and eigenvalues $\lambda_{j}(\omega)$.
+  3. iii
+    Derive dynamic principal component filters $\phi_{k}^{(j)}$.
+  4. iv
+    Calculate dynamic principal component scores $\text{dpc}_{t}^{(j)}$.
+3. Partial Out Influence of Other Nodes:
+$Y_{t}=Y^{*}_{t}-\hat{E}\bigl{(}Y^{*}_{t}\big{|}\mathbf{p}_{t}\bigr{)}$
+$X_{t}=X^{*}_{t}-\hat{E}\bigl{(}X^{*}_{t}\big{|}\mathbf{p}_{t}\bigr{)}$
 4. Fit Restricted/Unrestricted Models:
-    Unrestricted Model:
-    $Y_{t}=\alpha_{0}+\sum_{i=1}^{p}\alpha_{i}Y_{t-i}+\sum_{j=1}^{q}\beta_{j}X_{t-j}+\epsilon_{t}$
-    Restricted Model:
-    $Y_{t}=\gamma_{0}+\sum_{i=1}^{p}\gamma_{i}Y_{t-i}+\eta_{t}$
-    5. Perform Granger Causality Tests:
   1. i
-      Test null hypothesis $H_{0}:\beta_{j}=0$
-    for all $j$ in the equation for $Y_{t}$.
-  • ii
-    Compute F-statistic:
-    $F=\frac{(RSS_{R}-RSS_{U})/q}{RSS_{U}/(T-K)}$
-    where:
-    • $RSS_{R}$: Residual sum of squares of
-  • restricted model (without $X_{t-j}$ terms).
-  • $RSS_{U}$: Residual sum of squares of
-  • unrestricted model.
-  • $q$: Number of restrictions (parameters set to zero). $K$: Total number of parameters in the unrestricted model. iii Compare F-statistic with critical value
-  • from $F$-distribution $F_{q,T-K}$.
-  • iv
-    Reject $H_{0}$ if $F$ exceeds critical value; conclude $X_{t}$ Granger-causes $Y_{t}$. v Repeat steps for testing causality from $Y_{t}$ to $X_{t}$.
+    $Y_{t}=\alpha_{0}+\sum_{i=1}^{p}\alpha_{i}Y_{t-i}+\sum_{j=1}^{q}\beta_{j}X_{t-j}+\epsilon_{t}$.
+  2. ii
+    $Y_{t}=\gamma_{0}+\sum_{i=1}^{p}\gamma_{i}Y_{t-i}+\eta_{t}$.
+5. Testing Granger Causality:
+  1. i
+    Test $H_{0}:\beta_{j}=0$ for all $j$.
+  2. ii
+    Compute F-statistic $F=\frac{(RSS_{R}-RSS_{U})/q}{RSS_{U}/(T-K)}$.
+  3. iii
+    Reject $H_{0}$ if $F$ exceeds the critical value $F_{q,T-K}$.
 Procedure 1 Identifying Causal Connectivity in HDN
-    III Simulation study
-    The aim of this simulation study is to evaluate the effectiveness of spectral dynamic principal component analysis (sDPCA) combined with Granger causality (GC) testing in identifying true causal relationships within a high-dimensional network of time series data, specifically focusing on EEG-like data.
-The study investigates how well the proposed method isolates the channels of interest (COI) from the influences of other channels and accurately assesses their causal interactions.
-We conduct two sets of simulations: one with linear influences and another with nonlinear influences.
-    III-A Linear Scheme
-    In the linear scheme of the simulation study, we generate a network of time series data with predefined causal relationships and linear influences from additional channels.
-    III-A1 Generating Initial Channels
-    We begin by generating $N=40$ initial time series channels, denoted by $\{Z_{i}(t)\}_{i=1}^{N}$, over time points $t=1,\ldots,T$.
-The channels are divided into $N/2=20$ causal pairs, where each pair consists of a ”cause” and an ”effect” channel.
-    For each cause channel $Z_{i}(t)$, where $i=1,\ldots,N/2$, we model it as a weighted sum of latent processes representing different EEG frequency bands plus noise:
-    $Z_{i}(t)=a_{i}\,\delta(t)+b_{i}\,\theta(t)+c_{i}\,\alpha(t)+d_{i}\,\beta(t)+e_{i}\,\gamma(t)+\varepsilon_{i}(t),$
-    where:
-    • $\delta(t)$, $\theta(t)$, $\alpha(t)$, $\beta(t)$, and $\gamma(t)$ are latent processes corresponding to the EEG frequency bands:
-      • Delta (0.50.50.50.54444 Hz): $\delta(t)$,
-      • Theta (44448888 Hz): $\theta(t)$,
-      • Alpha (888812121212 Hz): $\alpha(t)$,
-      • Beta (1212121230303030 Hz): $\beta(t)$,
-      • Gamma ($>$ $30$ Hz): $\gamma(t)$.
-      Each latent process is modeled as an autoregressive process of order 2 (AR(2)):
-      $\nu(t)=\phi_{1\nu}\nu(t-1)+\phi_{2\nu}\nu(t-2)+\eta_{\nu}(t),\quad\nu\in\{\delta,\theta,\alpha,\beta,\gamma\},$
-      where $\eta_{\nu}(t)$ is white noise. The AR coefficients can expressed in terms of $r_{\nu}$ and $\varrho_{\nu}$:
-      $\phi_{1\nu}=2r_{\nu}\cos(\varrho_{\nu}),$
-      $\phi_{2\nu}=-r_{\nu}^{2}.$
-      where $r_{\nu}$ is the damping factor, chosen close to $1$ (e.g., $10.8\leq r_{\nu}<$ 1) to ensure stationarity and control the bandwidth of the spectral peak.
-And, $\varrho_{\nu}$ is the angular frequency (in radians per sample), related to the desired frequency $f_{\nu}$ (in Hz) by:
-      $\varrho_{\nu}=2\pi\frac{f_{\nu}}{f_{s}},$
-      with $f_{s}$ being the sampling frequency. By modeling the latent processes to correspond to specific EEG frequency bands, we aim to simulate EEG-like signals more realistically.
-    • $a_{i}$, $b_{i}$, $c_{i}$, $d_{i}$, and $e_{i}$ are random weights drawn from a uniform distribution to ensure variability among channels.
-    • $\varepsilon_{i}(t)$ is white noise added to each channel.
-    For each effect channel $Z_{j}(t)$, where $j=i+N/2$ and $i=1,\ldots,N/2$, we generate it as an autoregressive process influenced by its corresponding cause channel:
-    $Z_{j}(t)=\sum_{k=1}^{p}\phi_{k}Z_{i}(t-k)+\varepsilon_{j}(t),$
-    where:
-    • $p=2$ is the order of the autoregressive process.
-    • $\phi_{k}$ are AR coefficients ensuring the stability of the process.
-    • $\varepsilon_{j}(t)$ is white noise added to the effect channel.
-    III-A2 Extending the Network
-    We extend the network by generating additional $N_{\text{ext}}=472$ channels, resulting in a total of $N_{\text{expanded}}=512$ channels.
-These additional channels $\{Z_{i}(t)\}_{i=N+1}^{N_{\text{expanded}}}$ are generated similarly to the cause channels but without being part of the predefined causal pairs:
-    $Z_{i}(t)=a_{i}\,\delta(t)+b_{i}\,\theta(t)+c_{i}\,\alpha(t)+d_{i}\,\beta(t)+e_{i}\,\gamma(t)+\varepsilon_{i}(t).$
-    III-A3 Introducing Linear Influences
-    To simulate the effect of a high-dimensional network on the initial channels, we introduce linear influences from a subset of the additional channels.
-We select a set of connected nodes $\mathcal{C}\subset\{N+1,\ldots,N_{\text{expanded}}\}$ of size nofcn (number of connected nodes).
-    For each initial channel $Z_{i}(t)$, $i=1,\ldots,N$, we randomly select $k$ influencing nodes from $\mathcal{C}$ and assign influence weights $W_{ij}$ drawn from a uniform distribution.
-    The linear influence on channel $i$ at time $t$ is defined as:
-    $L_{i}(t)=\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t),$
+3 Simulation study
+This section presents a framework for evaluating whether Granger causality analysis, combined with sDPCA-based removal of confounding effects, can recover underlying causal associations in a set of EEG-like time series networks.
+The main objective is to demonstrate how this approach addresses linear, nonlinear, and causative external influences in high-dimensional settings.
+The third scenario is evaluated in a comparative framework.
+In this context, GC analysis with conventional PCA, DFM-based approaches by Barigozzi et al.
+(2024), and sparse vector autoregressive (VAR) estimation through parameter penalization techniques by (Nicholson et al., 2017, 2020) can be summoned up to obtain GC associations in large networks.
+These methods are made available in the R environment.
+Although the DFM estimation provided by the fnets package and the penalized VAR estimations provided by the BigVAR package are not directly designed for GC analysis, they can be regarded as alternatives since the estimated sparse coefficient matrices can signify GC links among variables.
+We have also considered alternatives where tests for GC detection can be performed with these techniques.
+All simulations incorporate designed connections among specific channels, while outer nodes of networks introduce confounding influences to obscure them.
+Suppose the total size of the generated networks is $N+N_{external}$ in which signals are labeled as $X_{1},\ldots,X_{N/2},Y_{1},\ldots,Y_{N/2},Z_{1},\ldots,Z_{N_{\mathrm{ext}}}$.
+The goal is to explore how well one can detect the causal connection $X_{i}\to Y_{i}$ after these pairs are affected by additional channels $\{Z_{j}\}$.
+Each $X^{*}_{i}$ is produced by mixing latent oscillatory components that reflect distinct frequency bands commonly used to approximate EEG activity.
+These bands—delta (0.50.50.50.54444 Hz), theta (44448888 Hz), alpha (888812121212 Hz), beta (1212121230303030 Hz), and gamma ($>$ $30$ Hz)—are treated as AR(2) processes of the form $\nu(t)=\psi_{1\nu}\nu(t-1)+\psi_{2\nu}\nu(t-2)+\eta_{\nu}(t),$ where $\psi_{1\nu}=2r_{\nu}\cos(\varrho_{\nu})$, $\psi_{2\nu}=-r_{\nu}^{2}$, and $\eta_{\nu}(t)\sim\mathrm{WN}(0,\sigma^{2}_{\nu})$.
+The resulting latent components, for $\nu\in\{\delta,\theta,\alpha,\beta,\gamma\}$, allow each $\nu$ to maintain a stationary and nearly resonant character.
+A typical $X_{i}$ can be written as
+$X_{i}(t)=a_{i}\delta(t)+b_{i}\theta(t)+c_{i}\alpha(t)+d_{i}\beta(t)+e_{i}\gamma(t)+\varepsilon_{i}(t),$
+where $\{a_{i},b_{i},c_{i},d_{i},e_{i}\}$ are channel-specific weights and $\varepsilon_{i}\sim\mathrm{WN}(0,\sigma^{2}_{\varepsilon})$.
+Its paired $Y^{*}_{i}$ is constructed by applying an AR model to $X^{*}_{i}$:
+$Y_{i}(t)=\sum_{k=1}^{p}\psi_{k}X_{i}(t-k)+\xi_{i}(t),\text{ where }\xi_{i}(t)\sim\mathrm{WN}(0,\sigma^{2}_{\xi})$
+with $p=2$ and stationarity-providing coefficients $\psi_{k}\in\pm 1.5$ in this study.
+This pairing establishes underlying cause-and-effect association $X_{i}\to Y_{i}$.
+An extended network is formed by adding exterior channels $\{Z_{j}\}$ to the initial causal setting, each generated in a similar fashion to $X_{i}$ but not designed to have cause-effect connections with the original pairs.
+These outer signals form a large repository of possible confounders.
+The final set contains $N_{\text{expanded}}$ time series, with the first $N$ holding the $N/2$ cause-effect relationships and the additional $N_{\text{external}}$ acting as external influencing signals that can inject disturbances into the original structure.
+3.1 Linear Scheme
+In order to simulate the confounding effect of a high-dimensional network on the initial channels, linear influences are introduced from a subset of the additional channels $\{Z_{j}\}$ to blur the association between $X_{i}$ and $Y_{i}$.
+We select a set of connected nodes $\mathcal{I}\subset\{1,\ldots,N_{ext}\}$.
+For each initial channel $Z_{j}(t)$, $j=1,\ldots,N_{ext}$, we randomly select $k$ influencing nodes from $\mathcal{I}$ and assign influence weights $W_{ij}$ drawn from a uniform distribution.
+Then, the linear influence (i.e., $L_{i}(t)=\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t)$) on channel $i$ at time $t$ can be given as:
+${X}^{*}_{i}(t)=X_{i}(t)+\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t),$
+${Y}^{*}_{i}(t)=Y_{i}(t)+\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t),$
 where $\mathcal{I}_{i}$ is the set of influencing nodes for channel $i$.
-For each initial channel $Z_{i}(t)$ and each influencing node $Z_{j}(t)\in\mathcal{I}_{i}$, we assign an influence weight $W_{ij}$ drawn from a uniform distribution centered around a specified influence weight $\omega$:
-    $W_{ij}\sim\text{Uniform}\left(\omega-\frac{\omega}{16},\,\omega+\frac{\omega}{16}\right),\quad\text{for }j\in\mathcal{I}_{i}.$
-    We then update the initial channels by adding the linear influence:
-    $\tilde{Z}_{i}(t)=Z_{i}(t)+L_{i}(t).$
-    The updated initial channels $\tilde{Z}_{i}(t)$ now contain both their original dynamics and the linear influence from the connected nodes.
-    III-B Nonlinear Scheme
-    In the nonlinear scheme, we introduce moderate nonlinearity into the influence of the connected nodes on the initial channels to study the robustness of the sDPCA and GC testing framework under nonlinear conditions.
-    III-B1 Introducing Nonlinear Influences
-    Similar to the linear simulation, we select a set of connected nodes $\mathcal{C}$ and assign influence weights $W_{ij}$ for each initial channel $Z_{i}(t)$.
-    The linear influence is computed as before:
-    $L_{i}(t)=\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t).$
-    To introduce nonlinearity, we compute interaction terms between the influencing nodes for each initial channel:
-    $Q_{i}(t)=\sum_{j\in\mathcal{I}_{i}}\sum_{k>j}W_{ij}W_{ik}Z_{j}(t)Z_{k}(t).$
-    We apply a nonlinear function $f(\cdot)$ to the linear influence, such as the softplus function:
-    $f(x)=\ln(1+e^{x}).$
-    The total influence on channel $i$ at time $t$ is then given by:
-    $S_{i}(t)=f\left(L_{i}(t)\right)+\omega Q_{i}(t),$
-    where $\omega$ is a scaling parameter controlling the magnitude of the nonlinear interaction influence.
-    We update the initial channels with the total influence:
-    $\tilde{Z}_{i}(t)=Z_{i}(t)+S_{i}(t).$
-    III-C Removing Influences
-    III-C1 Removing Influences Using sDPCA
-    To isolate the channels of interest from the influences of the connected nodes, we apply sDPCA to the time series data of the connected nodes $\{Z_{j}(t)\}_{j\in\mathcal{C}}$.
-We obtain principal component scores $\{Z_{\text{pc},k}(t)\}_{k=1}^{N_{\text{dpc}}}$, where $N_{\text{dpc}}$ is the number of dynamic principal components retained.
-For each adjusted initial channel $\tilde{Z}_{i}(t)$, we regress it on the principal component scores and their interaction terms:
-    $\tilde{Z}_{i}(t)=\sum_{k=1}^{N_{\text{dpc}}}\varphi_{ik}Z_{\text{pc},k}(t)+\sum_{k\neq l}\psi_{ikl}Z_{\text{pc},k}(t)Z_{\text{pc},l}(t)+\epsilon_{i}(t).$
-    We then use the residuals $\hat{\epsilon}_{i}(t)$ as the adjusted time series for further analysis:
-    $\hat{Z}_{i}(t)=\hat{\epsilon}_{i}(t).$
-    III-C2 Granger Causality Testing
-    Using the adjusted time series $\{\hat{Z}_{i}(t)\}_{i=1}^{N}$, we fit vector autoregressive (VAR) models and perform Granger causality tests to identify causal relationships among the initial channels.
-The effectiveness of the sDPCA adjustment is evaluated by comparing the detected causal relationships with the ground truth.
-    III-D Simulation Results
-    To evaluate the performance of the proposed method, we compute accuracy metrics based on the detection of true causal relationships.
-Specifically, we compare the identified causal links with the known ground truth and compute measures such as accuracy rate.
-    By conducting both linear and nonlinear simulations, we assess the robustness of the sDPCA combined with Granger causality testing in high-dimensional settings with complex interdependencies.
-The results provide insights into the method’s effectiveness in isolating the channels of interest and accurately detecting causal relationships, even in the presence of nonlinear influences.
-    The results are sensitive to the choice of latent processes and the influence of other channels.
-Different scenarios are explored by varying the number of influencing channels and the weight of their influence, with the goal of assessing the robustness and reliability of the proposed method in detecting true causal relationships within high-dimensional time series networks.
-For example, Table I presents the accuracy results for one of the simulation studies for different numbers of influencing nodes and varying weights on the COI in a high-dimensional network.
-The COI consists of the first 40 nodes out of a total of 512 nodes.
-The results of the other 5 simulation studies are given in the appendix.
-    TABLE I: Accuracy results for varying numbers of influencing nodes with nonlinear influence and weights on the COI in a high-dimensional network under a non-normal error distribution, $\epsilon\sim\mathrm{non\text{-}Norm}(\mu=-1,\sigma=1,\gamma=1.5,\kappa=3.5).$ The COI consists of the first 40 nodes out of a total of 512 nodes.
-The numbers 30, 60, 90, and 120 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, and 0.7 represent the weights of this influence.
-    Influencing Nodes = 30
-    Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-    3 0.96 0.88 0.61 0.65
-    6 0.97 0.77 0.62 0.64
-    12 0.97 0.82 0.61 0.65
-    24 0.93 0.82 0.65 0.64
-    Influencing Nodes = 60
-    Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-    3 0.98 0.66 0.64 0.64
-    6 0.99 0.73 0.64 0.65
-    12 0.94 0.69 0.64 0.64
-    24 0.96 0.67 0.64 0.64
-    Influencing Nodes = 90
-    Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-    3 0.96 0.63 0.63 0.65
-    6 0.94 0.62 0.64 0.64
-    12 0.95 0.62 0.64 0.64
-    24 0.91 0.64 0.64 0.64
-    Influencing Nodes = 120
-    Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
-    3 0.95 0.65 0.63 0.64
-    6 0.95 0.65 0.64 0.65
-    12 0.90 0.64 0.65 0.64
-    24 0.97 0.64 0.64 0.64
-    This simulation evaluates the performance of the sDPCA method under a nonlinear influence scheme with non-normal error terms, introducing complexity due to both nonlinearity and deviations from normality.
-    At a low influence weight (weight = 0.1), sDPCA maintains high accuracy across all numbers of influencing nodes, even as the network becomes super connected and super dense.
-Accuracy exceeds 0.90 when the number of influencing nodes increases from 30 up to 120, indicating strong robustness in isolating the channels of interest (COI) even in highly interconnected networks where each node affects the causal pairs.
-    As the influence weight increases to 0.3, accuracy decreases but remains relatively high with fewer influencing nodes.
-For example, with 30 influencing nodes and Ndpc = 3, the accuracy is 0.88.
-However, as the network becomes denser (influencing nodes ¿ 30), the decline in accuracy becomes more pronounced, reflecting the increased challenge of isolating true causal relationships in a super connected network.
-    At higher influence weights (weight = 0.5 and 0.7), accuracy declines further and stabilizes around 0.64–0.65, regardless of the number of influencing nodes or the number of dynamic principal components (Ndpc).
-This plateau suggests a limit to the method’s ability to isolate the COI under strong nonlinear influences and pervasive connectivity, as heavier weights and a denser network make it more challenging to distinguish true causal relationships from spurious ones introduced by nonlinearity and widespread influence.
-    The impact of increasing Ndpc on accuracy is minimal at higher weights and in denser networks, indicating that adding more principal components does not enhance performance under these conditions.
-This suggests that the sDPCA method’s capacity to capture and remove complex influences is limited when the network is both highly connected and subject to strong nonlinear interactions.
-    Overall, the sDPCA method demonstrates robustness to nonlinear influences and non-normal error distributions at lower influence weights and in less densely connected networks.
-Its effectiveness diminishes as the strength of the nonlinear influence increases and as the network becomes super connected, highlighting the importance of considering both the magnitude of external influences and network connectivity when applying sDPCA in practical settings with nonlinear dynamics.
-    IV Application to EEG data
-    In this section, we apply our proposed approach to real EEG data from an experiment involving motor execution and imagery tasks, offering insights into the temporal dynamics of brain activity.
-EEG recordings were collected from 109 volunteers using the BCI2000 system with a 64-channel setup; the data is publicly available at PhysioNet.org [18].
+For each initial channel $X^{*}_{i}(t)$ and each influencing node $Z_{j}(t)\in\mathcal{I}_{i}$, we assign an influence weight $W_{ij}$ drawn from a uniform distribution centered around a specified influence weight $\omega$ ( i.e., $W_{ij}\sim\text{Uniform}(\omega-\frac{\omega}{16},\omega+\frac{\omega}{16})$).
+The modified signals ${X}^{*}_{i}$ and ${Y}^{*}_{i}$ retain their causal arrangement in principle, but it is obscured by superimposed signals from the outer set.
+In the case of large amplitude weights (i.e., higher connectivity), the underlying causal association will be difficult to detect unless these background effects are adequately addressed.
+3.2 Nonlinear Scheme
+In the nonlinear scheme, we introduce moderate nonlinearity into the initial channels.
+The outer channels $\{Z_{j}\}$ remain potential confounders, but each channel ${X}^{*}_{i}$ or ${Y}^{*}_{i}$ includes products $Z_{j}(t)Z_{k}(t)$ and a smooth function of linear influence $L_{i}(t)=\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t)$.
+The nonlinear influence on channel $i$ at time $t$ can be exemplified as:
+${X}^{*}_{i}(t)=X_{i}(t)+f\Bigl{(}\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t)\Bigr{)}+\omega^{*}\sum_{j\in\mathcal{I}_{i}}\sum_{k>j}W_{ij}W_{ik}Z_{j}(t)Z_{k}(t),$
+${Y}^{*}_{i}(t)=Y_{i}(t)+f\Bigl{(}\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t)\Bigr{)}+\omega^{*}\sum_{j\in\mathcal{I}_{i}}\sum_{k>j}W_{ij}W_{ik}Z_{j}(t)Z_{k}(t),$
+where $f(x)=\ln\bigl{(}1+e^{x}\bigr{)}$ and $\omega^{*}$ scales the interaction term.
+3.3 Causal Scheme
+A separate simulation setup imposes cause-effect links among initial nodes and uses a broader set of $\{Z_{j}\}$ to create a complex background.
+The outer channels modify each $X_{i}$ or $Y_{i}$ according to difference equations like
+${X}^{*}_{i}(t)=\mu X_{i}(t-1)+\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t-1)+\nu_{i}(t),\text{ where }\nu_{i}(t)\sim\mathrm{WN}(0,\sigma^{2}_{\nu})$
+${Y}^{*}_{i}(t)=\mu Y_{i}(t-1)+\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t-1)+\nu_{i}(t),\text{ where }\nu_{i}(t)\sim\mathrm{WN}(0,\sigma^{2}_{\nu})$
+making it difficult to recover the direct link $X_{i}\to Y_{i}$ unless one accounts for external influences. This form is used as a framework in which multiple strategies can be compared.
+3.4 Simulation Results
+Linear and Nonlinear Schemes of Influences
+We assessed the proposed approach’s ability to recover known causal connections in simulated high-dimensional networks having both linear and nonlinear interference from outer channels explained in 3.1 and 3.2.
+In these settings, different numbers of influencing nodes (with varying weights) on the COI in a high-dimensional network are introduced.
+The size of the networks is set to 512 nodes.
+Thus, the COI consists of the first 40 (i.e., $N=40$) nodes out of a total of 512 (i.e., $N_{expanded}=512$) nodes.
+Figure 1 summarizes the results across these multiple setups.
+Results show that the procedure generally maintains high accuracy when confounding is modest.
+As the influence weight and network density grow, the method’s performance declines and stabilizes, indicating that strong external perturbations and extensive connectivity cause significant challenges.
+Increasing the number of dynamic principal components beyond a small set does not substantially improve accuracy under heavier confounding.
+However, the number of scores can be chosen according to their ability for the total variance explainability.
+In these scenarios, we focused more on a relatively small number of scores rather than explainability, so, total variance explainability changes between $20\%$ to $65\%$.
+Overall, the results suggest that spectral dynamic principal component analysis can effectively isolate the channels of interest in large networks, though very high connectivity or influence weights reduce its success.
+Figure 1: Accuracy results under Linear and Nonlinear Influences
+The method achieved high accuracy (above 0.90) when influence weights were low (e.g., 0.1), even as the number of interfering nodes increased from 30 to 120.
+As the influence weight rose to 0.3, performance remained acceptable for smaller interfering sets but declined with increasing network density.
+When the influence weights reached 0.5 and 0.7, the accuracy obtained was around 0.64–0.65, unaffected by the number of dynamic principal components used.
+This suggests that under strong nonlinear disturbances and widespread or super connectivity, the method’s ability to disentangle the channels of interest becomes limited.
+In lower interference regimes, sDPCA consistently filtered out background effects and preserved meaningful causal relationships.
+However, in highly entangled networks with amplified nonlinear effects, performance deteriorates regardless of the dimensionality reduction depth.
+These findings point to the method’s strengths in moderately complex environments and its limitations in ultra-dense or strongly nonlinear settings.
+Causative Schemes of Influences
+In another simulation scenario, we consider the case where the outer network acts in a causative manner given in 3.3.
+The size of the networks is set to 100 nodes in this case.
+The COI consists of the first 20 (i.e., $N=20$) nodes out of a total of 100 (i.e., $N_{expanded}=100$) nodes.
+Figure 2 shows boxplots of the Matthews correlation coefficient (MCC) and Cohen’s Kappa in the upper panels, along with line plots of mean values in the lower panels.
+Each plot illustrates how well different methods detect the injected causal links under varying degrees of network density (low, moderate, and high connectivity).
+The horizontal axis indicates the number of influencing nodes, which grows from low to high levels of confounding.
+GC–PCA uses standard principal component analysis to remove shared background signals before testing for Granger causality.
+DFM1, the default procedure, extracts idiosyncratic components via dynamic factor modeling and fits a sparse VAR to interpret nonzero coefficients as causal links.
+DFM2 modifies this by applying a formal Granger test to the idiosyncratic signals, rather than relying solely on a coefficient matrix.
+BigVAR1 fits an elastic-net penalized VAR to the influenced network, while BigVAR2 uses the same penalization strategy but only on factor-based idiosyncratic components extracted, forming a hybrid of DFM and penalized VAR estimation.
+Figure 2: Comparison of accuracies under Causative Influences
+Granger causality combined with sDPCA tends to provide higher accuracy scores across all scenarios.
+GC-PCA also performs well, though becomes less noticeable when the network density increases.
+Unlike the previous scenario, here, the number of scores used in partialling out of the effect of the outer network was determined when the cumulative explained variance reached $75\%$.
+Fitting BigVAR to only the idiosyncratic components estimated via DFM (BigVAR2) improves performance relative to fitting it to the full network (BigVAR1).
+DFM1 (default mode) shows relatively low performance in these simulations, whereas DFM2 (a conventional test for GC on the idiosyncratic components) recovers more causal links but remains below the top-performing procedures.
+Accuracy measures decrease as the number of influencing nodes grows, reflecting the increased difficulty of recovering true causal connections in super-connected systems.
+Methods that factor out or model the common background, such as sDPCA or the hybrid approach in BigVAR2, tend to be more robust in this regard.
+4 Application to EEG data
+In this section, we apply our proposed approach to EEG recordings from an experiment involving motor execution and imagery tasks, offering insights into the temporal dynamics of brain activity.
+EEG signals were collected from multiple subjects using the BCI2000 system with a 64-channel setup; the data is publicly available at PhysioNet.org (Schalk et al., 2004).
 Each subject completed 12 task-oriented runs.
-Participants either physically performed or imagined movements in response to visual cues on a screen, such as opening and closing fists or feet depending on the target’s position.
+Participants either physically performed or imagined movements in response to visual cues on a screen, such as opening and closing fists or feet, depending on the target’s position.
 The tasks were divided into four categories: physical movements of the left or right fist, imaginary movements of the left or right fist, physical movements of both fists or both feet, and imaginary movements of both fists or both feet.
@@
 All data underwent pre-processing to filter out artifacts like eye blinks and muscle movements, preserving the temporal structure of the recordings.
-The procedure included a high-pass filter at 0.5 Hz to eliminate low-frequency noise, removal of 60 Hz AC line noise, and Independent Component Analysis (ICA) using MNE-Python routines [19, 20] to isolate and exclude artifact components.
-Finally, artifacts associated with eye blinking and muscle movements were identified and removed, ensuring the recordings were free from these specific interferences.
+The procedure included a high-pass filter at 0.5 Hz to eliminate low-frequency noise, removal of 60 Hz AC line noise, and Independent Component Analysis (ICA) using MNE-Python routines (Gramfort et al., 2013; Larson et al., 2024) to isolate and exclude artifact components.
 Despite standard pre-processing, some recordings had unresolved issues like sawtooth traces, excessive blinking, broken channels, and unbalanced epochs.
 Therefore, we limited our analysis to 35 right-handed subjects (17 females and 18 males) whose recordings were qualified.
-To demonstrate how the proposed method can dissect causal connections between pairwise nodes in any time series network, we focused on six channels of interest (COI): AF3 and AF4 positioned over the pre-frontal cortex, C3 and C4 placed over the motor cortex, and O1 and O2 over the occipital cortex, as illustrated in Figure 1.
+To demonstrate how the proposed method can dissect causal connections between pairwise nodes in a network, we focused on six channels of interest (COI) out of 64 EEG channels: AF3 and AF4 positioned over the pre-frontal cortex, C3 and C4 placed over the motor cortex, and O1 and O2 over the occipital cortex, as illustrated in Figure 3(a).
 These channels were chosen because they correspond to brain regions likely to be active during voluntary movement, imagery, and visual perception.
@@
 This helps us to assess the efficacy of the proposed methodology for isolating COI while maintaining the temporal characteristics of the signals.
-    In Figure 1, we illustrate an example of one epoch from the dataset.
-The obtained three principal component (PC) scores are plotted over the time series data from which they were derived.
-This illustrates that the PC scores can capture the temporal structures/dynamics of the high-dimensional time series from which they originate and preserve the main characteristics of the time series.
-    Figure 1: Isolating Nodes via Spectral Dynamic PCA method
-    We compared the time-dependent statistical properties before and after applying the spectral dynamic principal component analysis (PCA) method to the channels of interest.
-As illustrated in Figures 2, 3 and 4, although the temporal characteristics of the series maintained, there is a noticeable trend toward stationarity in the serial correlation amplitudes, which showed signs of non-stationarity before.
-This adjustment indicates that while the sequential properties of the series are preserved, they become more aligned with the assumptions required by the auxiliary model used for Granger causality (GC) testing.
+In Figure 3(b), we illustrate an example of one epoch from the dataset.
+Three sDPCA scores obtained from excluded nodes are plotted over the time series data from which they were derived.
+This illustrates that the small number of scores can capture the global modes and preserve temporal structures/dynamics of the time series.
+Figure 3: (a)  : COI,  : Excluded nodes (b) sDPCA scores from Excluded nodes
+We compared the time-dependent properties before and after applying sDPCA to the channels of interest.
+As illustrated in Figures 4, 5 and 6, although the temporal characteristics of autocorrelation are maintained, there is a noticeable trend toward less correlatedness at long lags, suggesting that the shared variance with other brain areas has been removed.
+This can be considered as an indication of isolation from influencing nodes.
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
 Figure 6: Sample vs Isolated channels O1 and O2.
-    Test results given in Table II, reveal significant differences in GC tests conducted before and after the isolation process.
-In several instances, the GC tests showed contrasting results after the isolation (i.e., dimension reduction), underlining the presence of spurious GC regarding 64-dimensional VAR-based GC analysis.
-We observed that some of the causal relationships detected in the pre-isolation (i.e., 64-dimensional VAR) were not present in the post-isolation (i.e., 6-dimensional VAR).
-This suggests that the influences of other channels, when not addressed properly, can lead to misleading conclusions about causality.
-On the other hand, estimating causal relationships in networks with many time series —using VAR models and Granger Causality tests- is quite challenging, both statistically and computationally.
-The main issue is the ”curse of dimensionality.” As the number of variables increases, the number of parameters we need to estimate grows rapidly, leading to a paralyzing parameter space.
-This can result in overfitting, higher variance in estimators, and problems like multicollinearity when variables are highly correlated.
-Conventional estimation methods like Ordinary Least Squares (OLS) become unfeasible in this context.
-Additionally, selecting the right lag order and variables is more complicated with so many variables, increasing the risk of misspecification and inflated type I errors due to multiple testing.
-On the computational side, high-dimensional VAR models require significant resources that burst computational time when dealing with multiple epochs and subjects.
-This makes the Granger Causality testing process less reliable and efficient in these high-dimensional settings.
-    TABLE II: Comparison of GC Before and After Applying sDPCA
-    Direction of GC GC test w.r.t 64-dim VAR (Before) GC test w.r.t 6-dim VAR (After)
-    1C3 $\rightarrow$ C4 $\blacksquare$ $\blacksquare$
-    1C3 $\rightarrow$ AF3 $\blacksquare$ $\square$
-    1C3 $\rightarrow$ AF4 $\square$ $\blacksquare$
-    1C3 $\rightarrow$ O1 $\blacksquare$ $\blacksquare$
-    1C3 $\rightarrow$ O2 $\blacksquare$ $\square$
-    1C4 $\rightarrow$ C3 $\square$ $\blacksquare$
-    1C4 $\rightarrow$ AF3 $\square$ $\square$
-    1C4 $\rightarrow$ AF4 $\square$ $\blacksquare$
-    1C4 $\rightarrow$ O1 $\square$ $\square$
-    1C4 $\rightarrow$ O2 $\square$ $\blacksquare$
-    AF3 $\rightarrow$ AF4 $\square$ $\blacksquare$
-    AF3 $\rightarrow$ C3 $\square$ $\square$
-    AF3 $\rightarrow$ C4 $\square$ $\square$
-    AF3 $\rightarrow$ O1 $\square$ $\square$
-    AF3 $\rightarrow$ O2 $\square$ $\blacksquare$
-    AF4 $\rightarrow$ AF3 $\square$ $\blacksquare$
-    AF4 $\rightarrow$ C3 $\square$ $\square$
-    AF4 $\rightarrow$ C4 $\blacksquare$ $\blacksquare$
-    AF4 $\rightarrow$ O1 $\square$ $\square$
-    AF4 $\rightarrow$ O2 $\square$ $\square$
-    FO1 $\rightarrow$ O2 $\square$ $\blacksquare$
-    FO1 $\rightarrow$ AF3 $\blacksquare$ $\square$
-    FO1 $\rightarrow$ AF4 $\square$ $\square$
-    FO1 $\rightarrow$ C3 $\blacksquare$ $\blacksquare$
-    FO1 $\rightarrow$ C4 $\blacksquare$ $\square$
-    FO2 $\rightarrow$ O1 $\square$ $\blacksquare$
-    FO2 $\rightarrow$ AF3 $\square$ $\blacksquare$
-    FO2 $\rightarrow$ AF4 $\square$ $\blacksquare$
-    FO2 $\rightarrow$ C3 $\square$ $\blacksquare$
-    FO2 $\rightarrow$ C4 $\square$ $\square$
-    **Black squares, $\blacksquare$, indicate a ”Cause”, while white squares, $\square$, represent a ”Do not Cause” relationship.
-**Black squares, $\blacksquare$, indicate a ”Cause”, while white squares, $\square$, represent a ”Do not Cause” relationship.
-    In this section, we also estimate and present the connectivities for COI in the context of data from 35 subjects pertaining to both motor execution (ME) and motor imagery (MI) tasks.
+The remainder of this section presents estimated connectivities among COI regarding 35 subjects pertaining to both motor execution (ME) and motor imagery (MI) tasks.
 The subjects performed a series of opening and closing (OC) gestures, including left-hand fist (LH), right-hand fist (RH), both fists’ (BH), and both feet’ (BF), for both the ME and MI tasks.
@@
 From these results, the connectivities shared by at least 70% of the subjects were determined and demonstrated in Figure 7.
-    (a) Connectivity at Left Fist MI
-    (b) Connectivity at Left Fist ME
-    (c) Connectivity at Right Fist MI
-    (d) Connectivity at Right Fist ME
-    (e) Connectivity at Both Fists MI
-    (f) Connectivity at Both Fists ME
-    (g) Connectivity at Both Feet MI
-    (h) Connectivity at Both Feet ME
-    Figure 5: Connectivity at MI vs Connectivity at ME.
+(a) Left Fist (MI)
+(b) Right Fist (MI)
+(c) Both Fists (MI)
+(d) Both Feet (MI)
+(e) Left Fist (ME)
+(f) Right Fist (ME)
+(g) Both Fists (ME)
+(h) Both Feet (ME)
+Figure 7: Connectivity under Motor Imagery (top row) vs. Motor Execution (bottom row). Each panel (a)–(h) shows connectivity for Left Fist, Right Fist, Both Fists, and Both Feet.
 The GC results revealed distinct patterns of connectivity among the six channels of interest during different motor tasks, as illustrated in Figure 7.
@@
 During execution, there was increased causality from AF3 and AF4 (pre-frontal cortices) to motor and occipital regions, suggesting greater involvement of executive functions and planning during actual movement.
-For both feet tasks, the connectivity was less pronounced, with fewer significant Granger causalities observed during both imagery and execution.
-This may imply that foot movements engage different neural pathways or that the linear assessment is insufficient to capture the complexity.
+For both feet tasks, the connectivity patterns involve interactions among motor, pre-frontal, and occipital channels in both imagery and execution.
+Imagery and execution paths share influence from C4 onto multiple channels, indicating a common mechanism connecting the right motor cortex to pre-frontal and occipital regions.
+Findings suggest that BF movements share overlapping causal networks in both imagery and execution, but the directionality and intensity slightly differ.
 Overall, the contrasts between motor imagery and execution reveal that actual movement involves more extensive and complex connectivity, particularly with increased involvement of pre-frontal regions during execution.
@@
 By efficiently isolating the nodes of interest and removing the confounding effects of other channels, one may focus on specific neural interactions without the computational burden of analyzing the entire high-dimensional network.
-    V Discussion and Conclusion
-    Our study aimed to enhance the accuracy of Granger causality analysis in high-dimensional time-series networks, focusing on EEG datasets.
-By using the sDPCA method, we effectively isolated the channels of interest from the influence of other channels in the network.
-Results from both simulation studies and real EEG signal analysis demonstrate the simplicity and effectiveness of our proposed method.
-    We found that the method maintained high accuracy with a moderate number of influencing nodes.
-However, its accuracy decreased as the number of influencing nodes increased or when the network approached super connectivity, like swarm-like neuronal connections.
-This highlights the sensitivity of Granger causality to external influences and emphasizes the need to accurately model such complexities in real-world data.
-    Applying the proposed approach to EEG signal networks provided empirical validation of its utility.
-The principal component scores derived from the spectral domain accurately captured the temporal dynamics of the high-dimensional signals.
-After applying sDPCA, we observed a convergence towards stationarity in the serial correlation amplitudes, bringing the data closer to the assumptions required for reliable Granger causality testing.
-Comparing the results before and after isolation revealed spurious causality identifications in the pre-isolation data, underlining the importance of removing extraneous influences for accurate causality assessment.
-    Despite these promising findings, there are some limitations to our study.
-The simulation framework, while comprehensive, may not capture all nuances of real EEG data.
-Additionally, the test results can be significantly affected by the choice of hyperparameters, such as the number of principal components, the weights, and the number of influencing nodes.
-These parameters need to be carefully selected based on the specific characteristics of the datasets under study.
-Furthermore, our method relies on linearity-based PCA in the frequency domain, assuming that the underlying temporal dynamics can be adequately captured, which may not always hold true.
-Future research should consider integrating sDPCA with approaches that capture nonlinear relationships.
-To further improve the robustness and accuracy of causal inference in complex networks, exploring other dimensionality reduction techniques and advanced Granger causality methods would be beneficial.
-Moreover, the applicability of the methodology could be enhanced by including more sophisticated models of external influences and by incorporating real-time analysis capabilities.
-    Overall, our proposed approach offers a practical solution for isolating and analyzing causal associations in high-dimensional time series networks.
-Its ability to accurately identify causal connections while mitigating the effects of confounding influences makes it a valuable tool for researchers in neuroscience and other fields dealing with complex signals.
-The encouraging results of the study suggest that this method can improve the reliability of causality analysis, leading to more accurate and insightful investigations of the causal dynamics in high-dimensional networks.
-    VI Acknowledgments
-    We acknowledge the assistance of Grammarly and DeepL AI tools for grammar checks and language enhancement during the preparation of this manuscript.
+5 Discussion and Conclusion
+Our study aimed to enhance the usage of conventional Granger causality analysis in high-dimensional time-series networks, including application on EEG datasets.
+To this end, we investigated whether a two‑step strategy -sDPCA for dimensionality reduction followed by tests for Granger causality- can recover causal links between selected time series within a high-dimensional network.
+By projecting the activity of hundreds of background channels onto a handful of dynamic principal scores, the procedure lowers the parameter count, shortens computation time, and maintains the temporal information essential for EEG-like brain monitoring data.
+Simulated data confirmed that the proposed approach keeps false discoveries low and detects underlying connections if the background influence is not overwhelmingly strong.
+The EEG study provided a secondary line of evidence: after removing global confounding activity with sDPCA, the remaining channel pairs showed causal patterns corresponding to the known physiology of motor tasks.
+The results suggest that the approach captures spatial and temporal structure well enough to dissect meaningful interactions while avoiding the cost of fitting a full high‑dimensional VAR.
+Several conclusions emerge.
+First, spectral decomposition supplies a compact set of scores that retain lag and frequency-based information under low-dimensional representation, allowing conventional F‑tests to remain valid.
+Second, the reduction step adds little overhead; its computational burden scales mainly with the Fourier transform used to build the spectral matrix, not the square of the channel count.
+Third, the isolation step mitigates the long lag serial correlation of signals, which helps to uncover connections between series more efficiently.
+Together, these points make the workflow practical when the goal is to focus on one or two lead-lag relations embedded in a large network.
+The same findings also mark the limitations of the framework.
+When many channels share very strong drivers, relevant variance spreads across more dynamic components than can be kept without restoring noise, and detection power falls.
+Results also depend on a stable spectral density; abrupt shifts in frequency content undermine the eigen‑decomposition.
+Standard sDPCA relies on symmetric filters that are unusable in our context; replacing them with one‑sided causal filters prevents leakage of future information at the cost of lower frequency resolution.
+Finally, researchers interested in global connectivity still need sparse multivariate models.
+The proposed approach is not a universal replacement for full network modeling.
+When the scientific aim extends from a pair of nodes to the global network topology, sparse high‑dimensional modeling strategies remain necessary.
+Future extensions could embed the current sDPCA setting within richer pipelines.
+One path is to follow the causal scores with sparse‑ or ridge‑regularised VAR estimation, while another is to expand the score vector with kernel or other nonlinear transformations so that interaction terms and threshold effects enter the GC testing.
+A sequential variant could let the one‑sided filter update online, allowing real‑time monitoring.
+The dimensionality‑reduction stage can be upgraded by forming a weighted blend of latent factors drawn from nonlinear ICA, autoencoder embeddings, and graph‑spectral filters; this composite factor set feeds into the same purge‑then‑GC workflow and equips the method to address a wider spectrum of confounding patterns.
+In summary, sDPCA, followed by Granger causality testing, offers a practical compromise between full‑network modeling and bivariate analyses.
+It delivers reproducible inference on targeted node pairs while controlling for high‑dimensional background activity, provided that stationarity and moderate sparsity hold.
+These qualities make the approach useful for the toolkit for studying directed interactions in large‑scale time‑series data, such as EEG.

### v4

@@
 Such a representation is constructed through spectral dynamic principal component analysis (sDPCA) (Brillinger, 1969, 1981), which extends the principal component analysis to incorporate temporal dependence.
-By working with the spectral density matrix —directly linked to all auto and cross-covariances- sDPCA preserves frequency-specific features critical for neural data analysis.
+By working with the spectral density matrix, directly linked to all auto and cross-covariances, sDPCA preserves frequency-specific features critical for neural data analysis.
 The sDPCA enhances the PCA by incorporating temporal dependency within the data, thereby reducing the dimensionality of time series in the frequency domain while retaining temporal information.
@@
 Since there is a 1-1 relationship between the spectral matrix (at all frequencies) and the covariance matrix (at all lags), this approach captures lagged cross-correlations amongst signals.
+From a Bayesian paradigm, the generalized probabilistic PCA developed by Gu and Shen (2020) provides a dimension-reduction mechanism that accounts for serial and spatial correlation in multivariate signals.
 Various methodologies have been derived from frequency domain PCA.
@@
 Dynamic factor models (DFMs) (Forni et al., 2000, 2015; Stock and Watson, 2005; Barigozzi and Hallin, 2017) similarly rely on a few latent components, focusing on decomposing each series into common and idiosyncratic components.
-DFM allows more detailed lagged structures than static factor approaches but often seeks to explain broad data variance or perform large-scale forecasts.
+DFM allows more detailed lagged structures than static factor approaches, but often seeks to explain broad data variance or perform large-scale forecasts.
+Bayesian mixture network models can likewise account for cross-graph heterogeneity (Yin et al., 2022).
 Conditional Granger causality (CGC) (Geweke, 1984) instead tests whether adding one predictor refines forecasts of another, holding all others fixed.
@@
 The proposed approach distinguishes itself from existing approaches by focusing on frequency-domain dimensionality reduction combined with targeted pairwise causality analysis.
-The theoretical formulations of conditional GC, FAVAR, and DFMs have different aims – identifying influences through conditioning, incorporating large information sets into VARs, and explaining common variance in high dimensions, respectively – and each comes with certain assumptions about the data-generating process.
+The theoretical formulations of CGC, FAVAR, and DFMs have different aims – identifying influences through conditioning, incorporating large information sets into VARs, and explaining common variance in high dimensions, respectively – and each comes with certain assumptions about the data-generating process.
 The proposed approach does not compete with these state-of-the-art methods on their own terms; rather, it occupies a complementary cavity.
@@
 Empirically, this translates to a convenient method that is sensitive to the features of brain monitoring data – handling many channels and frequency-specific processes – while building on the proven strengths of factor modeling and Granger causality.
-Given the objectives of established methods and our approach, it is important to note that utilizing GC with sDPCA does not propose a universal substitute for full network modeling.
+Given the objectives of our approach, it is important to note that GC combined with sDPCA does not propose a universal substitute for full network modeling.
 Instead, it is a domain-orientated alternative that emphasizes interpretability and feasibility in the analysis of high-dimensional network time series.